### v2

# Compatibility of Kazhdan and Brauer Homomorphism

Abstract.
Let $G$ be a split connected reductive group, defined over $\mathbb{Z}$. Let $F$ and $F^{\prime}$ be the non-archimedean $m$-close local fields, where $m$ is a positive integer. D.Kazhdan gave an isomorphism between the Hecke algebras ${\rm Kaz}_{m}^{F}:\mathcal{H}\big{(}G(F),K_{F}\big{)}\rightarrow\mathcal{H}\big{(}G(F^{\prime}),K_{F^{\prime}}\big{)}$, where $K_{F}$ and $K_{F^{\prime}}$ are the $m$-th usual congruence subgroups of $G(F)$ and $G(F^{\prime})$ respectively. On the other hand, if $\sigma$ is an automorphism of $G$ of prime order $l$, then we have Brauer homomorphism ${\rm Br}:\mathcal{H}(G(F),U(F))\rightarrow\mathcal{H}(G^{\sigma}(F),U^{\sigma}(F))$, where $U(F)$ and $U^{\sigma}(F)$ are compact open subgroups of $G(F)$ and $G^{\sigma}(F)$ respectively. This article aims to study the compatibility between these two maps in certain sense. Further, an application of this compatibility is given in the context of linkage–which is the representation theoretic version of Brauer homomorphism.
1. Introduction
The local Langlands correspondence relates the set of irreducible smooth representations of the group of rational points of a reductive group over a local field–with the representation theory of its Weil–Deligne group. For a split reductive group G defined over $\mathbb{Z}$, Kazhdan conjectured a link between the local Langlands correspondences for $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$, where $F$ is a $p$-adic field and $F^{\prime}$ is a non-Archimedean equal characteristic local field which are sufficiently close. Kazhdan’s approach is via isomorphisms between the respective Hecke algebras, with integral coefficients, $\mathcal{H}(\operatorname{{\bf G}}(F),K)$ and $\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K^{\prime})$, for some specific choice of open-compact subgroups $K$ and $K^{\prime}$ of $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$ respectively (see [Kaz86, Theorem A]). Say $L$ is a non-Archimedean local field with residue characteristic $p$ and assume that $K$ is a nice compact open subgroup of $\operatorname{{\bf G}}(L)$–for instance a congruence subgroup of positive level. When $\operatorname{{\bf G}}(L)$ has an automorphism, denoted by $\sigma$, of prime order $l\neq p$ such that $\sigma(K)=K$, Treumann–Venkatesh ([TV16, Section 4]) define an $\overline{\mathbb{F}_{l}}$-algebra homomorphism between $\mathcal{H}(\operatorname{{\bf G}}(L),K)\otimes\overline{\mathbb{F}_{l}}$ and $\mathcal{H}(\operatorname{{\bf G}}^{\sigma}(L),K^{\sigma})\otimes\overline{\mathbb{F}_{l}}$, which is called the Brauer homomorphism. The Brauer homomorphism is conjectured to be compatible with Langlands functoriality principle. In this article, we study the compatibility of Brauer homomorhism and Kazhdan homomorphism in the Base change setting. Let $\mathfrak{o}_{L}$ be the ring of integers of $L$ and let $\mathfrak{p}_{L}$ be the maximal ideal of $\mathfrak{o}_{L}$. Let $E$ be a finite extension of $L$ of prime degree $l$, and let $\sigma$ be the generator of ${\rm Gal}(E/L)$. Let $K_{E}$ be the $m$-th congruence subgroup of $\operatorname{{\bf G}}(\mathfrak{o}_{E})$. Note that $K_{E}$ is stable under the action of $\sigma$ and let $K_{L}=K_{E}\cap\operatorname{{\bf G}}(L)$. Truemann and Venkatesh define Brauer homomorphism ${\rm Br}$, which is the algebra homomorphism
${\rm Br}:\mathcal{H}(\operatorname{{\bf G}}(E),K_{E})^{\sigma}\rightarrow\mathcal{H}(\operatorname{{\bf G}}(L),K_{L})$
obtained by restriction of functions on $\operatorname{{\bf G}}(E)$ to $\operatorname{{\bf G}}(L)$(see [TV16, Section 4]). The homomorphism ${\rm Br}$ is conjectured to be compatible with Langlands functoriality principle. Kazhdan ([Kaz86, Theorem A]) showed that there exists an integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close (see Section (4)), there is a Hecke algebra isomorphism
${\rm Kaz}_{m}^{F}:\mathcal{H}(\operatorname{{\bf G}}(F),K_{F})\rightarrow\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K_{F^{\prime}}).$
Kazhdan conjectured that it is sufficient to take $r=m$, and this conjecture is proved by Ganapathy in [Gan15]. It is shown in [Gan15] that Kazhdan isomorphism is compatible with the various properties in representation theory, and it also establish a bridge between local Langlands correspondence(LLC) over the field characteristic zero and that of positive characteristic.
We now set up some notations to formulate the main result of this article. Let $E$ be a finite Galois extension of a non-archimedean local field $F$ of prime degree $l$. Let $F^{\prime}$ be another non-Archimedean local field which is $m$-close to $F$. Throughout this article, we assume that the prime number $l$ is different from $p$, the residue characteristic of both $F$ and $F^{\prime}$. Lemma 4.2 in Section 4 says that we have a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ such that the fields $E$ and $E^{\prime}$ are $em$-close, where $e=1$ if $E$ is unramified and $e=l$ if $E$ is totally ramified. Let $\sigma$ (resp. $\sigma^{\prime}$) be a generator of the cyclic group ${\rm Gal}(E/F)$ (resp. ${\rm Gal}(E^{\prime}/F^{\prime})$). This leads to the following diagram:
[…]
In this article, we address the natural question about the commutativity of the above diagram. Let us state it as a theorem.
Theorem 1.1.
Let $F$ and $F^{\prime}$ be the non-archimedean $m$-close local fields. Let $E$ be the finite Galois extensions of $F$ of prime degree $l$, and let $E^{\prime}$ be the Galois extension of $F^{\prime}$, as indicated above. Then for any split connected reductive group G defined over $\mathbb{Z}$, we have
${\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E}={\rm Kaz}_{m}^{F}\circ{\rm Br}.$
In the context of representation theory, there is a notion of linkage (see subsection 6.2, for definition), defined by Treumann-Venkatesh–which is the representation theoretic analouge of Brauer homomorphism in the sense that they are compatible ([TV16, Section 6]). In this article, we give a precise formulation (subsection 6.3) of the compatibility of Brauer homomorphism with linkage. Then, we use this formulation and Theorem 1.1 to establish the compatibility of linkage with Kazhdan isomorphism (Theorem 6.3).
Now we briefly explain the contents of this article. In Section $2$, we review the Hecke algebra structure associated to any locally profinite group. In Section $3$, we formulate the Brauer homomorphism in our setting. The Kazdhan isomorphism and its compatibility with Galois actions will be discussed in Section $4$. In Section $5$, we prove Theorem 1.1. In Section $6$, we recall the notion of linkage and formulate some finiteness results on Tate cohomology group. Following this, we discuss the compatibility of linkage with Kazhdan isomorphism.
2. Hecke Algebra
In this section, we review the Hecke algebra structure on any locally profinite group. We refer to [TV16, Subsection 2.10] for more insight.
2.1.
Let $k$ be an algebraically closed field of positive characteristic $l$. Let $G$ be a locally profinite (i.e., locally compact and totally disconnected) group. Let $K$ be a compact open subgroup of $G$. We denote by $G/K$ the set of all distinct left cosets of $K$ in $G$, and it is a discrete set with a left action of $G$. For any $g\in G$, the left coset $gK$ is sometimes denoted by the symbol $[g]$. There is a natural action of $G\times G$, via left translation, on the space of all $k$-valued functions on the discrete set $G/K\times G/K$. We denote by $\mathcal{F}(G//K)$ the space of all such $k$-valued functions on $G/K\times G/K$ that are invariant under the action of the diagonal subgroup $\Delta G=\{(g,g):g\in G\}$, and whose support is a union of finitely many $G$-orbits. There is a multiplicative structure on the space $\mathcal{F}(G//K)$, given by
$(f_{1}*f_{2})([x],[z])=\sum_{y\in G/K}f_{1}([x],[y])f_{2}([y],[z]),$
for all $[x],[z]\in G/K$ and $f_{1},f_{2}\in\mathcal{F}(G//K)$. The space $\mathcal{F}(G//K)$ with the above multiplicative structure is a $k$-algebra, and it is called the Hecke algebra of $G$ associated with $K$. We now give an equivalent description of the Hecke algebra, which is relevant for our context.
2.2.
Let us consider the equivalence relation $\sim$ on the set $G/K\times G/K$, where
$([x],[y])\sim([x^{\prime}],[y^{\prime}])$
if and only if there exists $g\in G$ such that $[gx]=[gx^{\prime}]$ and $[gy]=[gy^{\prime}]$. Then for any $[g],[h]\in G/K$, the map $([g],[h])\mapsto Kg^{-1}hK$ induces a bijection from the set of equivalence classes $(G/K\times G/K)/\sim$ onto the double coset space $K\setminus G/K$.
Let $\mathcal{H}(G,K)$ be the space of all $k$-valued functions which are bi-$K$-invariant i.e., $f(k_{1}gk_{2})=f(g),$ for all $k_{1},k_{2}\in K$, $g\in G$ and $f\in\mathcal{H}(G,K)$. Then the preceding bijection gives the following map
$\mathcal{F}(G//K)\longrightarrow\mathcal{H}(G,K)$
$\phi\longmapsto\widetilde{\phi},$ (2.1)
where the function $\widetilde{\phi}:G\rightarrow k$ is defined as $\widetilde{\phi}(g):=\phi([1],[g])$, for all $g\in G$. The map (2.1) is clearly a bijection, and this induces a $k$-algebra structure on the space $\mathcal{H}(G,K)$. In the literature, many authors refer to the $k$-algebra $\mathcal{H}(G,K)$ as the Hecke algebra of $G$ associated with $K$.
3. Brauer Homomorphism
3.1.
Let $G$ be a locally profinite group, and let $\sigma$ be an automorphism of $G$ of order $l$. Let $K$ be a compact open subgroup of $G$ with $\sigma(K)=K$. Let $G^{\sigma}$ (resp. $K^{\sigma}$) be the the fixed subgroup of $G$ (resp. $K$) under the action of $\sigma$. Moreover, the compact open subgroup $K$ is called $\sigma$-plain if
  1. (1)
    The inclusion
    $G^{\sigma}/K^{\sigma}\hookrightarrow(G/K)^{\sigma}$
    $gK^{\sigma}\mapsto gK$
    is a bijection.
  2. (2)
    There exists a subgroup $K^{\prime}$ of $K$ of finite index such that $K^{\prime}$ is an inverse limit of finite subgroups each of which is coprime to $l$.
3.2.
For any $\sigma$-plain compact open subgroup $K$ of $G$, the discrete set $G/K\times G/K$ is equipped with an action of the group $\langle\sigma\rangle$, defined as
$\sigma.([g],[h]):=\big{(}[\sigma(g)],[\sigma(h)]\big{)},$
for all $[g],[h]\in G/K$. This action further factorizes through the quotient space $G/K\times G/K/\sim$, and it is given by
$\sigma.\overline{([g],[h])}=\overline{([\sigma(g)],[\sigma(h)])},$
where $\overline{([g],[h])}$ denotes the equivalence class of the element $([g],[h])$. The space $\mathcal{F}(G//K)$ carries a natural action of $\langle\sigma\rangle$, defined by
$(\sigma.\phi)([x],[y])=\phi\big{(}[\sigma^{-1}(x)],[\sigma^{-1}(y)]\big{)},$
for all $x,y\in G$ and $\phi\in\mathcal{F}(G//K)$. It is shown in [TV16, Section 4] that the Brauer homomorphism ${\rm Br}$ is a $k$-algebra homomorphism from $\mathcal{F}(G//K)^{\sigma}$ to $\mathcal{F}(G^{\sigma}//K^{\sigma})$, and it is the restriction map
${\rm Br}(\phi):={\rm Res}_{(G^{\sigma}/K^{\sigma}\times G^{\sigma}/K^{\sigma})}(\phi)$ (3.1)
for all $\phi\in\mathcal{F}(G//K)^{\sigma}$.
3.3.
Now, we formulate the Brauer homomorphism (3.1) for the $k$-algebra $\mathcal{H}(G,K)$, defined as in the preceding section. First, let us consider the action of the group $\langle\sigma\rangle$ on the double coset space $K\setminus G/K$ :
$\sigma.(KgK):=K\sigma(g)K,$
for all $g\in G$. Then we have a natural induced action of $\langle\sigma\rangle$ on the space $\mathcal{H}(G,K)$, defined as
$(\sigma.f)(g):=f(\sigma^{-1}(g)),$
for all $g\in G$. Note that the bijection (2.1) preserves the action of the group $\langle\sigma\rangle$. Then we have commutative diagram
${\mathcal{F}(G//K)^{\sigma}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm Br}}$${\mathcal{H}(G,K)^{\sigma}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm\overline{Br}}}$${\mathcal{F}(G^{\sigma}//K^{\sigma})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{H}(G^{\sigma},K^{\sigma})}$
where the horizontal arrows in the diagram are the bijections induced by (2.1). We observe that the $k$-algebra map ${\rm\overline{Br}}:\mathcal{H}(G,K)^{\sigma}\rightarrow\mathcal{H}(G^{\sigma},K^{\sigma})$, induced by the above diagram, is the restriction map to subgroup $G^{\sigma}$ i.e.,
${\rm\overline{Br}}(f):={\rm Res}_{G^{\sigma}}(f),$
for all $f\in\mathcal{H}(G,K)^{\sigma}$
4. Kazhdan isomorphism
In this section, we review Kazhdan isomorphism and set up some initial results. Let $F$ and $F^{\prime}$ be the non-Archimedean local fields, both of residue characteristic $p$, with ring of integers $\mathfrak{o}_{F}$ and $\mathfrak{o}_{F^{\prime}}$ and its maximal ideals $\mathfrak{p}_{F}$ and $\mathfrak{p}_{F^{\prime}}$ respectively. The fields $F$ and $F^{\prime}$ are called $m$-close if there is a ring isomorphism $\Lambda:\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\rightarrow\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$. We choose uniformizer $\varpi$ (resp. $\varpi^{\prime}$) of the field $F$ (resp. $F^{\prime}$) such that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the map $\Lambda$.
4.1.
Let G be a split connected reductive group defined over $\mathbb{Z}$. Fix a Borel subgroup B of G, defined over $\mathbb{Z}$. Let $\textbf{B}=\textbf{T}\textbf{U}$ be a Levi decomposition over $\mathbb{Z}$ with maximal torus T and unipotent radical U. We denote by $X^{*}(\textbf{T})$ and $X_{*}(\textbf{T})$ the character lattice and the cocharacter lattice of T respectively. Then there is a natural $\mathbb{Z}$-bilinear pairing
$X^{*}(\textbf{T})\times X_{*}(\textbf{T})\longrightarrow\mathbb{Z}$
$(\alpha,\lambda)\longmapsto\langle\alpha,\lambda\rangle,$
where $(\lambda\circ\alpha)(x)=x^{\langle\alpha,\lambda\rangle}$, for all $x\in\mathbb{G}_{m}$. Let $\Phi\subseteq X^{*}(\textbf{T})$ be the set of roots of T in G, and let $\Phi^{+}$ be the set of positive roots of T in B. Let $G_{F}=\textbf{G}(F)$. Consider the subgroup
$K_{F}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F})\longrightarrow\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\big{)}.$
Then $K_{F}$ is a compact open subgroup of $G_{F}$. Fix a Haar measure $\mu_{F}$ on $G_{F}$ such that $\mu_{F}(K_{F})=1$. For $g\in G_{F}$, let $t_{g}=1_{K_{F}gK_{F}}$, the characteristic function on the double coset $K_{F}gK_{F}$. Then the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ is generated as a $k$-vector space by $\{t_{g}:g\in G_{F}\}$. We denote by $\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})$ the space generated by $\{t_{g}:g\in G_{F}\}$ over $\mathbb{Z}$. Set
$X_{*}(\textbf{T})^{-}=\{\lambda\in X_{*}(\textbf{T}):\langle\alpha,\lambda\rangle\leq 0,\alpha\in\Phi^{+}\}.$
We have the Cartan decomposition
$G_{F}=\coprod_{\mu\in X_{*}(\textbf{T})^{-}}\textbf{G}(\mathfrak{o}_{F})\varpi_{\mu}\textbf{G}(\mathfrak{o}_{F}),$ (4.1)
where $\varpi_{\mu}=\mu(\varpi)$. For each $\mu\in X_{*}(\textbf{T})^{-}$, the double coset $G_{F}(\mu):=\textbf{G}(\mathfrak{o}_{F})\varpi_{\mu}\textbf{G}(\mathfrak{o}_{F})$ is a homogeneous space under the action of the group $\textbf{G}(\mathfrak{o}_{F})\times\textbf{G}(\mathfrak{o}_{F})$, defined as
$(x,y).g:=xgy^{-1},$
for all $x,y\in\textbf{G}(\mathfrak{o}_{F})$ and $g\in G_{F}(\mu)$. Then the finite group $H_{m}:=\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\times\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})$ acts transitively on the set $\{K_{F}xK_{F}:x\in G_{F}(\mu)\}$. Let $\Gamma_{\mu}(F)$ be the stabilizer of $K_{F}\varpi_{\mu}K_{F}$ in $H_{m}$. Let $T_{\mu}(F)\subseteq\textbf{G}(\mathfrak{o}_{F})\times\textbf{G}(\mathfrak{o}_{F})$ be the set of representatives of the coset space $H_{m}/\Gamma_{\mu}(F)$ under the composition
$\textbf{G}(\mathfrak{o}_{F})\times\textbf{G}(\mathfrak{o}_{F})\xrightarrow{{\rm mod}-\mathfrak{p}_{F}^{m}}H_{m}\longrightarrow H_{m}/\Gamma_{\mu}(F)$
Then (4.1) can be re-written as
$G_{F}=\coprod_{\mu\in X_{*}(\textbf{T})^{-}}\coprod_{(b_{i},b_{j})\in T_{\mu}(F)}K_{F}b_{i}\varpi_{\mu}b_{j}^{-1}K_{F}.$ (4.2)
For each $\mu\in X_{*}(\textbf{T})^{-}$, the map $\Lambda$ induces a bijection $T_{\mu}(F)\rightarrow T_{\mu}(F^{\prime})$. Under this bijection, there is an isomorphism ([Kaz86, Section 1]) of $\mathbb{Z}$-modules:
$\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})\longrightarrow\mathcal{H}_{\mathbb{Z}}(G_{F^{\prime}},K_{F^{\prime}})$
$t_{a_{i}\varpi_{\lambda}a_{j}^{-1}}\longmapsto t_{a_{i}^{\prime}\varpi_{\lambda}^{\prime}a_{j}^{\prime-1}},$
and this induces a vector space isomorphism
${\rm Kaz}_{m}^{F}:\mathcal{H}(G_{F},K_{F})\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}}),$
where $\lambda\in X_{*}(\textbf{T})^{-}$, $(a_{i},a_{j})\in T_{\lambda}(F)$ and $(a_{i}^{\prime},a_{j}^{\prime})$ is the corresponding element in $T_{\lambda}(F^{\prime})$. For $F$ and $F^{\prime}$, sufficiently close fields, the following result is due to Kazhdan ([Kaz86, Theorem A]).
Proposition 4.1.
Given $m\geq 1$. There exists a positive integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close, the map ${\rm Kaz}_{m}^{F}$ defined above is an algebra isomorphism.
Kazhdan conjectured that one could take $r=m$ in Proposition 4.1, as later proved by R.Ganapathy in her paper [Gan15].
4.2.
Let $F$ and $F^{\prime}$ be the non-Archimedean $m$-close local fields with residue chracteristic $p$. In [Del84], Deligne proved that there is an isomorphism
${\rm Gal}(\overline{F}/F)/\mathcal{I}_{F}^{m}\xrightarrow{{\rm Del}_{m}}{\rm Gal}(\overline{F^{\prime}}/F^{\prime})/\mathcal{I}_{F^{\prime}}^{m},$
where $\overline{F}$ (resp. $\overline{F^{\prime}}$) is the seperable closure of $F$ (resp. $F^{\prime}$), and $\mathcal{I}_{F}^{m}$ (resp. $\mathcal{I}_{F^{\prime}}^{m}$) is the $m$-th higher ramification subgroup of the inertia group $\mathcal{I}_{F}$ (resp. $\mathcal{I}_{F^{\prime}}$). Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$. Following the description of the map ${\rm Del}_{m}$, as mentioned in [Gan15, Section 2.1], we have a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of same degree $l$. Moreover, the extension $E^{\prime}/F^{\prime}$ can be constructed so that $E$ and $E^{\prime}$ are likewise sufficiently close, as ensured by the next lemma.
Lemma 4.2.
Let $F$ and $F^{\prime}$ be the $m$-close non-Archimedean local fields, and let $E$ be the finite Galois extension of $F$ of prime degree $l$. Then there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ such that $E$ and $E^{\prime}$ are $em$-close, where $e\in\{1,l\}$.
Proof.
The proof is divided into two parts. In the first part, we consider the case when $E$ is unramified over $F$. The second part deals with totally ramified case. Note that the extension $E/F$ is tamely ramified as $l\neq p$. We denote by $A_{m}$ and $A_{m}^{\prime}$ the quotient rings $\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ and $\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$ respectively. Recall that the isomorphism $\Lambda:A_{m}\xrightarrow{\sim}A_{m}^{\prime}$ takes the class of $\varpi$ to that of $\varpi^{\prime}$.
4.2.1. Unramified case
Suppose $E/F$ is unramified. In this case, the field $E$ is of the form $F[T]/(f)$, where $f$ is a monic irreducible polynomial with coefficients in $\mathfrak{o}_{F}$ such that its ${\rm mod}-\mathfrak{p}_{F}$ reduction, say $\overline{f}$, is irreducible. Moreover, the quotient ring $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{m}$ is isomorphic to $A_{m}[T]/(\overline{f})$. Let $\overline{f}^{\prime}$ be the image of $\overline{f}$ under the ring isomorphism $\Lambda$. Then we have
$\mathfrak{o}_{E}/\mathfrak{p}_{E}^{m}\xrightarrow{\sim}A_{m}^{\prime}[T]/(\overline{f}^{\prime}).$
Let $f^{\prime}$ be an irreducible polynomial in $\mathfrak{o}_{F^{\prime}}[T]$ with its ${\rm mod}-\mathfrak{p}_{F}^{m}$ reduction $\overline{f}^{\prime}$. Let $E^{\prime}$ be the field extension $F^{\prime}[T]/(f^{\prime})$. Then $E^{\prime}$ is unramified over $F^{\prime}$ and there is a ring isomorphism
$\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{m}\simeq A_{m}^{\prime}[T]/(\overline{f}^{\prime}).$
Thus we have the ring isomorphism $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{m}\xrightarrow{\sim}\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{m}$, which implies that the fields $E$ and $E^{\prime}$ are $m$-close.
4.2.2. Totally ramified case
Suppose $E$ is totally ramified over $F$. Here the field $E$ is of the form $F[T]/(T^{l}-\varpi)$ and its ring of integers $\mathfrak{o}_{E}$ is given by $\mathfrak{o}_{F}[T]/(T^{l}-\varpi)$. Recall that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the isomorphism $\Lambda$. Now consider the field extension
$E^{\prime}:=F[T]/(T^{l}-\varpi^{\prime})$
Then $E^{\prime}$ is totally ramified over $F^{\prime}$ with the ring of integers $\mathfrak{o}_{E^{\prime}}=\mathfrak{o}_{F^{\prime}}[T]/(T^{l}-\varpi^{\prime})$. Let $\overline{\varpi}$ be the image of $\varpi$ under ${\rm mod}-\mathfrak{p}_{F}^{m}$. Then the natural surjection
$\mathfrak{o}_{F}[T]/(T^{l}-\varpi)\longrightarrow A_{m}[T]/(T^{l}-\overline{\varpi})$
factorzies through the ideal $\mathfrak{p}_{F}^{m}[T]/(T^{l}-\varpi)$, and it gives the ring isomorphism
$\mathfrak{o}_{E}/\mathfrak{p}_{E}^{lm}\xrightarrow{\sim}A_{m}[T]/(T^{l}-\overline{\varpi}).$
Similarly, there is a ring isomorphism
$\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{lm}\simeq A_{m}^{\prime}[T]/(T^{l}-\overline{\varpi^{\prime}}),$
where $\overline{\varpi^{\prime}}$ is the image of $\varpi^{\prime}$ under the ${\rm mod}-\mathfrak{p}_{F^{\prime}}^{m}$ reduction map. On the other hand, the map $\Lambda$ induces the ring isomorphism
$A_{m}[T]/(T^{l}-\overline{\varpi})\xrightarrow{\sim}A_{m}^{\prime}[T]/(T^{l}-\overline{\varpi^{\prime}}).$
This shows that the ring $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{lm}$ is isomorphic to $\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{lm}$, and hence the fields $E$ and $E^{\prime}$ are $lm$-close. ∎
4.2.3.
For each $e\in\{1,l\}$, we fix a ring isomorphism $\Pi:\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\rightarrow\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}$ as above, induced by $\Lambda$, so that the diagram
${\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Lambda}$${\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}}$
is commutative, where the vertical arrows are the natural inclusion maps. We choose uniformizer $\pi$ (resp. $\pi^{\prime}$) of the field $E$ (resp. $E^{\prime}$) such that the class of $\pi$ is mapped to the class of $\pi^{\prime}$ under the isomorphism $\Pi$. For any $\overline{a}\in\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ (or, $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}$), we write $\overline{a}^{\prime}$ for the image $\Lambda(\overline{a})$ (or, $\Pi(\overline{a})$), whenever no confusion arises.
4.3. Compatibility with Galois action
Here, we continue to work with the set up of the preceeding subsection. We fix a generator $\sigma$ of the group ${\rm Gal}(E/F)$. Then $\sigma$ can be regarded as an automorphism of the group $G_{E}$, via its natural action on $G_{E}$. Consider the compact open subgroup
$K_{E}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{E})\rightarrow\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})\big{)}.$
We observe that $K_{E}$ is $\sigma$-plain, and there is an induced action of $\langle\sigma\rangle$ on the Hecke algebra $\mathcal{H}(G_{E},K_{E})$, given by
$(\sigma.f)(x)=f(\sigma^{-1}x),$
for all $x\in G_{E}$ and $f\in\mathcal{H}(G_{E},K_{E})$. Similarly, there is an action of $\langle\sigma^{\prime}\rangle$ on the Hecke algebra $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$. We choose a group isomorphism $\iota:{\rm Gal}(E/F)\rightarrow{\rm Gal}(E^{\prime}/F^{\prime})$, sending $\sigma$ to a generator, say $\sigma^{\prime}$, that makes commutative the diagram
${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$$\scriptstyle{\sigma}$${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma^{\prime}}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}}$
where the horizontal maps are the ring isomorphisms induced by $\sigma:E\rightarrow E$ and $\sigma^{\prime}:E^{\prime}\rightarrow E^{\prime}$. To be more precise, for any $a\in\mathfrak{o}_{E}$, we have the relation
$\overline{\sigma(a)}^{\prime}=\sigma(\overline{a})^{\prime}=\sigma^{\prime}(\overline{a}^{\prime})=\overline{\sigma^{\prime}(a^{\prime})},$ (4.3)
where $a^{\prime}\in\mathfrak{o}_{E^{\prime}}$ is the preimage of $\overline{a}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. We end this section with the following proposition, which shows the compatibility of Kazhdan isomorphism with these Galois actions.
Proposition 4.3.
For all $f\in\mathcal{H}(G_{E},K_{E})$, we have
${\rm Kaz}_{em}^{E}(\sigma.f)=\sigma^{\prime}.{\rm Kaz}_{em}^{E}(f).$
Proof.
In view of the Cartan decomposition (4.2), it is enough to show that
${\rm Kaz}_{em}^{E}(\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}})=\sigma^{\prime}.{\rm Kaz}_{em}^{E}(t_{a_{i}\pi_{\mu}a_{j}^{-1}}),$
for all $\mu\in X_{*}(\textbf{T})^{-}$ and $(a_{i},a_{j})\in T_{\mu}(E)$. Note that
$\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}}=t_{\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})}.$
By Cartan decomposition (4.2), the double coset $K_{E}\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})K_{E}$ is of the form $K_{E}c_{i}\pi_{\mu}c_{j}^{-1}K_{E}$, where $(c_{i},c_{j})\in T_{\mu}(E)$. Then
${\rm Kaz}_{em}^{E}(\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}})={\rm Kaz}_{em}^{E}\big{(}t_{\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j})^{-1}}\big{)}={\rm Kaz}_{em}^{E}\big{(}t_{c_{i}\pi_{\mu}c_{j}^{-1}}\big{)}=t_{c_{i}^{\prime}\pi_{\mu}^{\prime}c_{j}^{\prime-1}}.$ (4.4)
Consider the Galois conjugate $\sigma(\pi)=\pi t=\pi\sigma(s)$, where $t,s\in\mathfrak{o}_{E}^{\times}$. Under the relation $\sigma(\mu(x))=\mu(\sigma(x))$, for $x\in E^{\times}$, we observe that
$K_{E}c_{i}\pi_{\mu}c_{j}^{-1}K_{E}$ $=K_{E}\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})K_{E}$
$=K_{E}\sigma(a_{i})\pi_{\mu}\sigma(s_{\mu}a_{j}^{-1})K_{E}$
$=K_{E}\sigma(a_{i})\pi_{\mu}\sigma(d_{j}^{-1})K_{E},$
where $s_{\mu}=\mu(s)$, and $d_{j}=a_{j}s_{\mu}^{-1}\in\textbf{G}(\mathfrak{o}_{E})$. Then $\big{(}\bar{c_{i}}^{-1}\overline{\sigma(a_{i})},\overline{\sigma(d_{j}^{-1})}\bar{c_{j}}\big{)}\in\Gamma_{\mu}(E)$, and using the bijection $\Gamma_{\mu}(E)\rightarrow\Gamma_{\mu}(E^{\prime})$, we get
$\big{(}\bar{c_{i}}^{\prime-1}\overline{\sigma(a_{i})}^{\prime},\overline{\sigma(d_{j}^{-1})}^{\prime}\bar{c_{j}}^{\prime}\big{)}\in\Gamma_{\mu}(E^{\prime}).$
Recall that the image of $d_{j}$ under the reduction map $\textbf{G}(\mathfrak{o}_{E})\rightarrow\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})$ is denoted by $\overline{d_{j}}$, and $\overline{d_{j}}^{\prime}$ is the image of $\overline{d_{j}}$ under the isomorphism $\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})\xrightarrow{\sim}\textbf{G}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em})$. Let $d_{j}^{\prime}\in\textbf{G}(\mathfrak{o}_{E^{\prime}})$ be a preimage of $\overline{d_{j}}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. Then the commutative diagram
${\mathbb{G}_{m}(\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$$\scriptstyle{\mu}$${\textbf{T}(\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathbb{G}_{m}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$${\textbf{T}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}})}$
gives the relation
$\overline{d_{j}^{\prime}}=\overline{a_{j}^{\prime}}\overline{s_{\mu}}^{\prime-1}=\overline{a_{j}^{\prime}}(\overline{s^{\prime}_{\mu}})^{-1},$
where $s^{\prime}\in\mathfrak{o}_{E^{\prime}}$ is a preimage of $\overline{s}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$. On the other hand, the relation $\sigma(\pi)=\pi\sigma(s)$ together with (4.3) implies that
$\overline{\sigma^{\prime}(\pi^{\prime})}=\sigma^{\prime}(\overline{\pi}^{\prime})=\overline{\pi^{\prime}}\overline{\sigma^{\prime}(s^{\prime})}.$
For any $\mu\in X_{*}(\textbf{T})^{-}$, the above equality gives
$\overline{(\sigma^{\prime}(\pi^{\prime}))_{\mu}}=\overline{(\pi^{\prime}\sigma^{\prime}(s^{\prime}))_{\mu}}.$
This implies that $(\sigma^{\prime}(\pi^{\prime}))_{\mu}^{-1}(\pi^{\prime}\sigma^{\prime}(s^{\prime}))_{\mu}\in K_{E^{\prime}}$. Now, using normality of $K_{E^{\prime}}$ in $\textbf{G}(\mathfrak{o}_{E^{\prime}})$ and the relation $\sigma(\mu(x))=\mu(\sigma(x))$, for $x\in E^{\times}$, we get
$K_{E^{\prime}}c_{i}^{\prime}\pi_{\mu}^{\prime}c_{j}^{\prime-1}K_{E^{\prime}}$ $=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\pi_{\mu}^{\prime}\sigma^{\prime}(d_{j}^{\prime-1})K_{E^{\prime}}$
$=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})K_{E^{\prime}}.$
This shows that
$t_{c_{i}^{\prime}\pi_{\mu}^{\prime}c_{j}^{\prime-1}}=t_{\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})}=\sigma^{\prime}.t_{a_{i}^{\prime}\pi_{\mu}^{\prime}a_{j}^{\prime-1}}.$
Finally, it follows from (4.4) and the above relation that
${\rm Kaz}_{em}^{E}(\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}})=\sigma^{\prime}.{\rm Kaz}_{em}^{E}(t_{a_{i}\pi_{\mu}a_{j}^{-1}}).$
5. Compatibility of Brauer homomorphism and Kazhdan isomorphism
In this section, we prove Theorem 1.1. Let us first recall the notations and terminologies from the prceeding sections.
5.1.
Fix a positive integer $m$. Let $F$ and $F^{\prime}$ be the non-Archimedean local fields such that $F$ is $m$-close to $F^{\prime}$. Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$, where $p$ is the residue characteristic of both $F$ and $F^{\prime}$. According to Lemma 4.2, there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ and $E^{\prime}$ is $em$-close to $E$, where $e\in\{1,l\}$.
Let G be a split connected reductive group defined over $\mathbb{Z}$. It follows from [KP23, Proposition 12.9.4] that the fixed point subgroup $K_{E}^{\sigma}$ (resp. $K_{E^{\prime}}^{\sigma^{\prime}}$) is the $m$-th congruence subgroup of $\textbf{G}(\mathfrak{o}_{F})$ (resp. $\textbf{G}(\mathfrak{o}_{F^{\prime}})$) i.e, $K_{E}^{\sigma}=K_{F}$ and $K_{E^{\prime}}^{\sigma^{\prime}}=K_{F^{\prime}}$, where
$K_{F}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F})\rightarrow\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\big{)}$
and
$K_{F^{\prime}}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F^{\prime}})\rightarrow\textbf{G}(\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m})\big{)}.$
Proposition 4.3 then gives an isomorphism between $\mathcal{H}(G_{E},K_{E})^{\sigma}$ and $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}$, induced by ${\rm Kaz}_{em}^{E}$. On the other hand, we have the Brauer homomorphism $\overline{\rm Br}$ (resp. $\overline{\rm Br}^{\prime}$) from the space $\mathcal{H}(G_{E},K_{E})^{\sigma}$ (resp. $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}$) to the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ (resp. $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$). With these set up, we aim to prove
Theorem 5.1.
Let $F,F^{\prime},E$ and $E^{\prime}$ be as above. Then for any split connected reductive group G defined over $\mathbb{Z}$, we have
${\rm Kaz}_{m}^{F}\circ\overline{\rm Br}=\overline{\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E}.$
Proof.
The proof is divided into two parts. In the first part, we consider the extension $E/F$ to be unramified. The second part deals with totally ramified case. To start with, let us recall that an arbitrary $h\in\mathcal{H}(G_{E},K_{E})^{\sigma}$ can be written as
$h=\sum_{\mu\in X_{*}(\textbf{T})^{-}}\sum_{(a_{i},a_{j})\in T_{\mu}(E)}\alpha_{\mu}(a_{i},a_{j})t_{a_{i}\pi_{\mu}a_{j}^{-1}},$
where each $\alpha_{\mu}(a_{i},a_{j})\in k$. Then
$({\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}})(h)=\sum_{(a_{i},a_{j})\in T_{\mu}(E)}\alpha_{\mu}(a_{i},a_{j})({\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}})(t_{a_{i}\pi_{\mu}a_{j}^{-1}})$ (5.1)
and
$({\rm\overline{Br}}^{\prime}\circ{\rm Kaz}_{em}^{E})(h)=\sum_{\mu\in X_{*}(\textbf{T})^{-}}\sum_{(a_{i},a_{j})\in T_{\mu}(E^{\prime})}\alpha_{\mu}(a_{i},a_{j}){\rm\overline{Br}}^{\prime}(t_{a_{i}^{\prime}\pi_{\mu}^{\prime}a_{j}^{\prime-1}}).$ (5.2)
For each $\mu\in X_{*}(\textbf{T})^{-}$ and $(a_{i},a_{j})\in T_{\mu}(E)$, we have ${\rm\overline{Br}}(t_{a_{i}\pi_{\mu}a_{j}^{-1}})=1_{G_{F}\cap K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}}$. Then Cartan decomposition (4.2) gives
$K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}\cap G_{F}=\coprod_{\lambda\in X_{*}(\textbf{T})^{-}}\coprod_{(b_{i},b_{j})\in T_{\lambda}(F)}K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}\cap K_{F}b_{i}\varpi_{\lambda}b_{j}^{-1}K_{F}.$ (5.3)
Similarly,
$K_{E^{\prime}}a_{i}^{\prime}\pi_{\mu}^{\prime}a_{j}^{\prime-1}K_{E^{\prime}}\cap G_{F^{\prime}}=\coprod_{\nu\in X_{*}(\textbf{T})^{-}}\coprod_{(d_{i}^{\prime},d_{j}^{\prime})\in T_{\nu}(F^{\prime})}K_{E^{\prime}}a_{i}^{\prime}\pi^{\prime}_{\mu}a_{j}^{\prime-1}K_{E^{\prime}}\cap K_{F^{\prime}}d_{i}^{\prime}\varpi^{\prime}_{\nu}d_{j}^{\prime-1}K_{F^{\prime}}.$ (5.4)
5.1.1. Unramified case
Suppose $E$ is unramified over $F$. Then the field $E^{\prime}$ is also unramified over $F^{\prime}$ and it is $m$-close to $E$ (Lemma 4.2). In this case, $\pi=\varpi$ and $\pi^{\prime}=\varpi^{\prime}$. For any $\lambda\in X_{*}(\textbf{T})^{-}$ and $(b_{i},b_{j})\in T_{\lambda}(F)$, it follows from the decompositions (5.3) and (5.4) that
$b_{i}\varpi_{\lambda}b_{j}^{-1}\in K_{E}a_{i}\varpi_{\mu}a_{j}^{-1}K_{E}$
if and only if
$\varpi_{\lambda}\in K_{E}b_{i}^{-1}a_{i}\varpi_{\mu}a_{j}^{-1}b_{j}K_{E},$
and this will happen only if $\lambda=\mu$. Then
$\varpi_{\mu}\in K_{E}b_{i}^{-1}a_{i}\varpi_{\mu}a_{j}^{-1}b_{j}K_{E}$
if and only if
$(b_{i}^{-1}a_{i},b_{j}^{-1}a_{j})\in\Gamma_{\mu}(E).$
Using the bijection $\Gamma_{\mu}(E)\rightarrow\Gamma_{\mu}(E^{\prime})$ $(\overline{a}\mapsto\overline{a}^{\prime})$, we have
$(\overline{b_{i}}^{-1}\overline{a_{i}},\overline{b_{j}}^{-1}\overline{a_{j}})\in\Gamma_{\mu}(E)$
if and only if
$(\overline{b_{i}}^{\prime-1}\overline{a_{i}}^{\prime},\overline{b_{j}}^{\prime-1}\overline{a_{j}}^{\prime})=(\overline{b_{i}^{\prime}}^{-1}\overline{a_{i}^{\prime}},\overline{b_{j}^{\prime}}^{-1}\overline{a_{j}^{\prime}})\in\Gamma_{\mu}(E^{\prime}),$
where $(b_{i}^{\prime},b_{j}^{\prime})\in T_{\mu}(F^{\prime})$. Therefore, we have
$b_{i}\varpi_{\mu}b_{j}^{-1}\in K_{E}a_{i}\varpi_{\mu}a_{j}^{-1}K_{E}\iff b_{i}^{\prime}\varpi_{\mu}^{\prime}b_{j}^{\prime-1}\in K_{E^{\prime}}a_{i}^{\prime}\varpi_{\mu}^{\prime}a_{j}^{\prime-1}K_{E^{\prime}}$ (5.5)
In view of (5.5), it follows from (5.1) and (5.2) that
$({\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}})(h)=({\rm\overline{Br}}^{\prime}\circ{\rm Kaz}_{m}^{E})(h).$ (5.6)
5.1.2. Totally Ramified case
Here, we assume that the extension $E/F$ is totally ramified. Then Lemma 4.2 ensures that the extension $E^{\prime}/F^{\prime}$ is also totally ramified and $E^{\prime}$ is $lm$-close to $E$, where $l=[E:F]=[E^{\prime}:F^{\prime}]$. Note that
$K_{E}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{E})\xrightarrow{{\rm mod}-\mathfrak{p}_{E}^{lm}}\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{lm})\big{)}$
and
$K_{F}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F})\xrightarrow{\text{mod}-\mathfrak{p}_{F}^{m}}\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\big{)}.$
Let $\xi\in X_{*}(\textbf{T})^{-}$ and $(c_{i},c_{j})\in T_{\xi}(F)$. In view of the decompositions (5.3) and (5.4), we have
$c_{i}\varpi_{\xi}c_{j}^{-1}\in K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}$
if and only if
$\varpi_{\xi}\in K_{E}c_{i}^{-1}a_{i}\pi_{\mu}a_{j}^{-1}c_{j}K_{E},$
and it is possible only if $\xi^{l}=\mu$. Therefore, it follows that
$\pi_{\xi^{l}}\in K_{E}c_{i}^{-1}a_{i}\pi_{\xi^{l}}a_{j}^{-1}c_{j}K_{E}$
if and only if
$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{\xi^{l}}(E).$
Under the bijection $\Gamma_{\xi^{l}}(E)\rightarrow\Gamma_{\xi^{l}}(E^{\prime})$, we have
$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{\xi^{l}}(E^{\prime})$
if and only if
$(\overline{c_{i}}^{\prime-1}\overline{a_{i}}^{\prime},\overline{c_{j}}^{\prime-1}\overline{a_{j}}^{\prime})=(\overline{c_{i}^{\prime}}^{-1}\overline{a_{i}^{\prime}},\overline{c_{j}^{\prime}}^{-1}\overline{a_{j}^{\prime}})\in\Gamma_{\xi^{l}}(E^{\prime}),$
where $(c_{i}^{\prime},c_{j}^{\prime})\in T_{\xi^{l}}(F^{\prime})$. Thus we have
$c_{i}\varpi_{\xi}c_{j}^{-1}\in K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}\iff c_{i}^{\prime}\varpi^{\prime}_{\xi}c_{j}^{\prime-1}\in K_{E^{\prime}}a_{i}^{\prime}\pi^{\prime}_{\mu}a_{j}^{\prime-1}K_{E^{\prime}}.$ (5.7)
From the relations (5.1), (5.2) and (5.7), we get
$({\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}})(h)=({\rm\overline{Br}}^{\prime}\circ{\rm Kaz}_{lm}^{E})(h).$
This completes the proof. ∎
6. Linkage over close local fields
In this section, an application of Theorem 5.1 is given in the context of linkage. To be more precise, we show that the linkage is compatible with Kazhdan isomorphism. To prove this, we first give a formulation of the compatibility of linkage with Brauer homomorphism, as described in [TV16, Section 6.2], and then we use Theorem 5.1 to get the required compatibility. To start with, we first recall the Tate cohomology and the concept of linkage. For a reference, see [TV16, Section 3 and 6].
6.1. Tate cohomology
Let $\Gamma$ be a cyclic group of order $n$ with generator $\gamma$. Let $M$ be an abelian group with an action of $\Gamma$. Let $T_{\gamma}$ be the automorphism of $M$ defined by
$T_{\gamma}(m)=\gamma.m$, for $\gamma\in\Gamma,m\in M$.
Let $N={\rm id}+T_{\gamma}+T_{\gamma^{2}}+\dots+T_{\gamma^{n-1}}$ be the norm operator. The Tate cohomology groups $\widehat{H}^{0}(M)$ and $\widehat{H}^{1}(M)$ are defined as
$\widehat{H}^{0}(M):={\rm ker}({\rm id}-T_{\gamma})/{\rm Im}(N)$,
$\widehat{H}^{1}(M):={\rm ker}(N)/{\rm Im}({\rm id}-T_{\gamma})$.
6.1.1. Tate Cohomology of sheaves on $\ell$-spaces
Let $X$ be an $\ell$-space (i.e., Hausdorff, locally compact and totally disconnected) with an action of a finite group $\langle\gamma\rangle$ of prime order $l$. Let $\mathcal{F}$ be a sheaf of $k$-modules on $X$. If $\mathcal{F}$ is $\gamma$-equivariant, then $\gamma$ can be regarded as a map of restricted sheaves $\mathcal{F}|_{X^{\gamma}}\rightarrow\mathcal{F}|_{X^{\gamma}}$ and the Tate cohomology is defined as
$\widehat{H}^{0}(\mathcal{F}|_{X^{\gamma}}):={\rm ker}(1-\gamma)/{\rm Im}(N)$,
$\widehat{H}^{1}(\mathcal{F}|_{X^{\gamma}}):={\rm ker}(N)/{\rm Im}(1-\gamma)$.
For each $i\in\{0,1\}$, the object $\widehat{H}^{i}(\mathcal{F}|_{X^{\gamma}})$ is a sheaf on $X^{\gamma}$.
6.2. Linkage
Let $F$ be a non-archimedean local field and let $G$ be the $F$-points of a connected reductive group defined over $F$. A representation $(\rho,V)$ of $G$ is called smooth if for every vector $v\in V$, the $G$-stabilizer of $v$ is an open subgroup of $G$. Moreover, such a representation $\rho$ is called admissible if for each compact open subgroup $K$ of $G$, the space of $K$-fixed vectors $V^{K}$ (sometimes denoted by $\rho^{K}$) is finite-dimensional over $k$. In this subsection, all the representations are assumed to be admissible and the representation spaces are defined over $k$. We denote by $\mathcal{R}(G)$ the category of smooth admissible $k$-representations of $G$.
6.2.1. Frobenius Twist
Let $(\rho,V)$ be a representation of $G$. Consider the vector space $V^{(l)}$, where the underlying additive group structure of $V^{(l)}$ is same as that of $V$ but the scalar action $\star$ on $V^{(l)}$ is given by
$c\star v=c^{\frac{1}{l}}.v$ , for all $c\in k,v\in V$.
Then the action of $G$ on $V$ induces a representation $\rho^{(l)}$ of $G$ on $V^{(l)}$. The representation $(\rho^{(l)},V^{(l)})$ is called the Frobenius twist of the representation $(\rho,V)$.
6.2.2.
Let $\sigma$ be an automorphism of $G$ of prime order $l$. An irreducible representation $(\Xi,W)$ of $G$ is called $\sigma$-fixed if $\Xi$ is isomorphic to $\Xi\circ\sigma$. For such representation $\Xi$, there is an unique action of $\sigma$ on the space $W$, which is compatible with the action of $\sigma$ on $G$ ([TV16, Proposition 6.1]). Then we have the Tate cohomology space $\widehat{H}^{i}(W)$, defined as $k$-representation of $G^{\sigma}$. We denote this representation by $\widehat{H}^{i}(\Xi)$.
Definition 6.1.
An irreducible representation $\rho$ of $G^{\sigma}$ is linked with the representation $\Xi$ if the Frobenius-twist $\rho^{(l)}$ occurs as a Jordan-Holder constituent of $\widehat{H}^{0}(\Xi)$ or $\widehat{H}^{1}(\Xi)$.
6.2.3.
It was conjectured by Venkatesh-Treumann ([TV16, Section 6.3]) that $\widehat{H}^{i}(\Xi)$ is of finite length as a representation of $G^{\sigma}$. We give a proof of this conjecture for the group ${\rm GL}_{n}$, under the following set up.
[…]
We start with a non-archimedean local field $L$. Let $P_{n}(L)$ and $U_{n}(L)$ be the subgroups of ${\rm GL}_{n}(L)$, given by
$P_{n}(L)=\Big{\{}\begin{pmatrix}A&B\\ 0&1\end{pmatrix}:A\in{\rm GL}_{n-1}(L),B\in L^{n-1}\Big{\}}$
and
$U_{n}(L)=\Big{\{}\begin{pmatrix}1&C\\ 0&1\end{pmatrix}:C\in L^{n-1}\Big{\}}$
respectively. We denote by $X_{L}$ the coset space $P_{n}(L)/U_{n}(L)$. Let $\theta_{n}$ be the character of $U_{n}(L)$, defined as
$\theta_{n}((u)_{i,j}):=\psi_{L}(u_{n-1,n})$, for $(u)_{i,j}\in U_{n}(L)$.
We have four fundamental functors:
$\Psi^{-}:\mathcal{R}(P_{n}(L))\rightarrow\mathcal{R}({\rm GL}_{n-1}(L)),\Psi^{+}:\mathcal{R}({\rm GL}_{n-1}(L))\rightarrow\mathcal{R}(P_{n}(L))$
$\Phi^{-}:\mathcal{R}(P_{n}(L))\rightarrow\mathcal{R}(P_{n-1}(L)),\Phi^{+}:\mathcal{R}(P_{n-1}(L))\rightarrow\mathcal{R}(P_{n}(L)),$
given by $\Psi^{-}=r_{U_{n}(L),{\rm id}}$, $\Phi^{-}=r_{U_{n}(L),\theta_{n}}$, $\Psi^{+}=i_{U_{n}(L),{\rm id}}$, and $\Phi^{+}=i_{U_{n}(L),\theta_{n}}$.
Let $\tau$ be a smooth representation of $P_{n}(L)$. The $m$-th derivative of $\tau$, denoted by $\tau^{(m)}$, is constructed as the representation $\Psi^{-}(\Phi^{-})^{m-1}(\tau)$ of ${\rm GL}_{n-m}(L)$. Moreover, we get a functorial filtration on $\tau$:
$0\subseteq\tau_{n}\subseteq\tau_{n-1}\subseteq\dots\subseteq\tau_{2}\subseteq\tau_{1}=\tau,$
where $\tau_{m}=(\Phi^{+})^{m-1}(\Phi^{-})^{m-1}(\tau)$ and $\tau_{m}/\tau_{m+1}=(\Phi^{+})^{m-1}(\Psi^{+})(\tau^{(m)})$. Using these, we now deduce:
Proposition 6.2.
Let $\Xi$ be an irreducible $\sigma$-fixed representation of ${\rm GL}_{n}(E)$. Let $T_{\sigma}:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism with $T_{\sigma}^{l}={\rm id}$. Then the Tate cohomology $\widehat{H}^{i}(\Xi)$ with respect to the operator $T_{\sigma}$ is a finite length representation of ${\rm GL}_{n}(F)$.
Proof.
We use induction on the integer $n$. Assume that the proposition is true for all positive integers less than $n$. Since $\psi_{E}(\sigma(x))=\psi_{E}(x)$, for $x\in E$, we get the isomorphism
$\Phi^{-}(\Xi\circ\sigma)\simeq\Phi^{-}(\Xi)\circ\sigma,$ (6.1)
as representation of $P_{n-1}(E)$. Similarly, for any representation $\tau\in\mathcal{R}(P_{n-1}(E))$, we have the $P_{n}(E)$-equivariant isomorphism
$\Phi^{+}(\tau\circ\sigma)\simeq\Phi^{+}(\tau)\circ\sigma.$ (6.2)
The map $T_{\sigma}$ together with (6.1) and (6.2) induces an isomorphism between $\Xi^{(m)}$ and $(\Xi^{(m)})\circ\sigma$, and also between the representations $\Xi_{m}$ and $\Xi_{m}\circ\sigma$ for $m\geq 1$. Using induction hypothesis, we get that the representation $\widehat{H}^{i}(\Xi^{(m)})$ is a finite length representation of ${\rm GL}_{n-m}(F)$. Since $P_{n}(E)$ is stable under the action of $\sigma$, for any $n\geq 1$, we get that there is a bijection between the sets $X_{E}^{\sigma}$ and $X_{F}$. Then according to [TV16, Proposition 3.3.1], we have
$\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})=(\Phi^{+})^{m-1}(\Psi^{+})\big{(}\widehat{H}^{i}(\Xi^{(m)})\big{)}.$
Hence, for all $m\geq 1$, the representation $\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})$ is finite length representation of $P_{n}(F)$. Now, using spectral sequence induced by canonical filtration of ${\rm Res}_{P_{n}(E)}(\Xi)$, we get that $\widehat{H}^{i}(\Xi)$ is finite length representation of $P_{n}(F)$, and this implies that $\widehat{H}^{i}(\Xi)$ is finite length representation of ${\rm GL}_{n}(F)$. ∎
It is reasonable to expect that Proposition 6.2 holds true for any general connected reductive algebraic group G over $F$. In the next section, we give a formulation of the compatibility of Brauer homomorphism with linkage, under the finite length assumption of the Tate cohomology $\widehat{H}^{i}(\Xi)$ for any irreducible $\sigma$-fixed representation $\Xi$ of $G=\textbf{G}(F)$. Then we show that this formulaton is compatible with Kazhdan isomorphism.
6.3. Compatibility of Brauer homomorphism with linkage
Let $\Xi$ be an irreducible $\sigma$-fixed representation of $G$, and let $T:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism such that $T^{l}={\rm id}$. For any $\sigma$-plain compact open subgroup $K$ of $G$, the inclusion map $\Xi^{K}\hookrightarrow\Xi$ is $\sigma$-invariant, and it induces a map $\widehat{H}^{i}(\Xi^{K})\rightarrow\widehat{H}^{i}(\Xi)$, taking values in the $\mathcal{H}(G^{\sigma},K^{\sigma})$-module $\widehat{H}^{i}(\Xi)^{K^{\sigma}}$. It is shown in [TV16, Section 6.2] that for all $h\in\mathcal{H}(G,K)^{\sigma}$, the Brauer homomorphism $\overline{\rm Br}$ induces the commutative diagram
${\widehat{H}^{i}(\Xi^{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{H}^{i}(h)}$${\widehat{H}^{i}(\Xi)^{K^{\sigma}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm\overline{Br}}(h)}$${\widehat{H}^{i}(\Xi^{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{i}(\Xi)^{K^{\sigma}}}$
and this induces the compatibility of Brauer homomorphism with linkage in the sense that, for a representation $\rho$ of $G^{\sigma}$ to be linked with $\Xi$, it is equivalent to say that there exists a composition factor $N$ of the $\mathcal{H}(G,K)^{\sigma}$-module $\widehat{H}^{i}(\Xi^{K})$ such that
$\mathcal{H}(G^{\sigma},K^{\sigma})\otimes_{\mathcal{H}(G,K)^{\sigma}}N\simeq(\rho^{(l)})^{K^{\sigma}}.$ (6.3)
6.3.1. Compatibility of Kazhdan isomorphism with linkage
We first recall the notations and set up under which we study this compatibility. Let $F$ and $F^{\prime}$ be the non-Archimedean $m$-close local fields with same residue characteristic $p$. Let $E$ be the finite Galois extension of $F$ of prime degree $l$ with $l\neq p$. By Lemma 4.2, there exists a finite Galois extension $E^{\prime}/F^{\prime}$ with same degree of extension $l$. As before, we fix the generators $\sigma$ and $\sigma^{\prime}$ of the Galois groups ${\rm Gal}(E/F)$ and ${\rm Gal}(E^{\prime}/F^{\prime})$ respectively.
Let G be a split connected reductive group defined over $\mathbb{Z}$. Let $G_{E}$ (resp. $G_{F}$) be the $E$ (resp. $F$)-points of G. Let $K_{E}$ and $K_{F}$ be the $m$-th congruence subgroups of $G_{E}$ and $G_{F}$ respectively (see Section 4). Then we have
${\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}}={\rm\overline{Br}^{\prime}}\circ{\rm Kaz}_{em}^{E},$ (6.4)
where $e\in\{1,l\}$. It is shown in [Gan15, Section 2.3] that there are bijections
$\big{\{}\rho_{E}\in{\rm Irr}(G_{E}):\rho_{E}^{K_{E}}\neq 0\big{\}}\longrightarrow\big{\{}\rho_{E^{\prime}}\in{\rm Irr}(G_{E^{\prime}}):\rho_{E^{\prime}}^{K_{E^{\prime}}}\neq 0\big{\}}$ (6.5)
and
$\big{\{}\rho_{F}\in{\rm Irr}(G_{F}):\rho_{F}^{K_{F}}\neq 0\big{\}}\longrightarrow\big{\{}\rho_{F^{\prime}}\in{\rm Irr}(G_{F^{\prime}}):\rho_{F^{\prime}}^{K_{F^{\prime}}}\neq 0\big{\}},$ (6.6)
induced by the Kazhdan isomorphisms ${\rm Kaz}_{em}^{E}$ and ${\rm Kaz}_{m}^{F}$ respectively. With these notations and terminologies, we now prove
Theorem 6.3.
Let $(\rho_{E},V_{E})$ be an irreducible $\sigma$-fixed representation of $G_{E}$ with $V_{E}^{K_{E}}\neq 0$, and let $(\rho_{F},V_{F})$ be an irreducible representation of $G_{F}$ with $V_{F}^{K_{F}}\neq 0$. Let $(\rho_{E^{\prime}},V_{E^{\prime}})$ and $(\rho_{F^{\prime}},V_{F^{\prime}})$ be the corresponding objects under the bijections (6.5) and (6.6). Then $\rho_{F}$ is linked with $\rho_{E}$ if and only if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$.
Proof.
The proof is divided into two parts. In the first part, we show that the irreducible representation $\rho_{E^{\prime}}$, obtained from $\rho_{E}$ under the bijection (6.5), is $\sigma^{\prime}$-fixed. The second part shows that the linkage between $\rho_{F}$ and $\rho_{E}$ implies the linkage between $\rho_{F^{\prime}}$ and $\rho_{E^{\prime}}$, and vice versa.
6.3.2.
First recall that the space $V_{E}^{K_{E}}$ is an irreducible $\mathcal{H}(G_{E},K_{E})$-module, and it can be regarded as $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$-module via Kazhdan isomorphism. Consider the $\mathcal{H}(G_{E^{\prime}})$ (For definition, see [Vig96, Chapter 1])-module
$N^{\prime}=\mathcal{H}(G_{E^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})}V_{E}^{K_{E}}.$
Note that the action of $\langle\sigma^{\prime}\rangle$ on $\mathcal{H}(G_{E^{\prime}})$ induces an action of $\langle\sigma^{\prime}\rangle$ on $N^{\prime}$, given by
$\sigma^{\prime}.(f^{\prime}\otimes v)=(\sigma^{\prime}.f^{\prime})\otimes v,$
for $f^{\prime}\in\mathcal{H}(G_{E^{\prime}})$ and $v\in V_{E}^{k_{E}}$. Following the construction of [Vig96, Section 4, Chapter 1], the $\mathcal{H}(G_{E^{\prime}})$-module $N^{\prime}$ gives the irreducible representation $\rho_{E^{\prime}}$, defined as
$\rho_{E^{\prime}}(g^{\prime})(x^{\prime})=1_{g^{\prime}K_{E^{\prime}}}*x^{\prime},$
for $g^{\prime}\in G_{E^{\prime}},x^{\prime}\in V_{E^{\prime}}$. Here, *** denotes the action of the Hecke algebra $\mathcal{H}(G_{E^{\prime}})$ on the space $V_{E^{\prime}}$ and $1_{g^{\prime}K_{E^{\prime}}}$ is the characteristic function of the left coset $g^{\prime}K_{E^{\prime}}$. Since $\rho_{E}$ is $\sigma$-fixed, there exists an isomorphism $T_{\sigma}:\rho_{E}\rightarrow\rho_{E}\circ\sigma$. Then the map $T_{\sigma^{\prime}}:N^{\prime}\rightarrow N^{\prime}$, defined by
$T_{\sigma^{\prime}}(f^{\prime}\otimes v):=(\sigma^{\prime}.f^{\prime})\otimes T_{\sigma}(v),$
induces an isomorphism $\rho_{E^{\prime}}\xrightarrow{\sim}\rho_{E^{\prime}}\circ\sigma^{\prime}$, and this shows that the representation $\rho_{E^{\prime}}$ is $\sigma^{\prime}$-fixed.
6.3.3.
Now, suppose that the representation $\rho_{F}$ is linked with $\rho_{E}$. Then the $\mathcal{H}(G_{F},K_{F})$-module $(V_{F}^{(l)})^{K_{F}}$ is a composition factor of $\widehat{H}^{i}(V_{E})^{K_{F}}$, for some $i\in\{0,1\}$. If we regard these as $\mathcal{H}(G_{E},K_{E})^{\sigma}$-modules via the Brauer homomorphism $\overline{\rm Br}$, then the formulation (6.3) gives a composition factor $M$ of $\widehat{H}^{i}(V_{E}^{K_{E}})$ such that
$\mathcal{H}(G_{F},K_{F})\otimes_{\mathcal{H}(G_{E},K_{E})^{\sigma}}M\simeq(V_{F}^{(l)})^{K_{F}}.$ (6.7)
Note that the isomorphism $V_{E}^{K_{E}}\rightarrow V_{E^{\prime}}^{K_{E^{\prime}}}$ of $\mathcal{H}(G_{E},K_{E})$-modules induces a $\mathcal{H}(G_{E},K_{E})^{\sigma}$-module isomorphism $\Phi:\widehat{H}^{i}(V_{E}^{K_{E}})\rightarrow\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, which is also a $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}$-module isomorphism via the Kazhdan map ${\rm Kaz}^{E}_{em}$.
Let $M^{\prime}$ be the image of $M$ under the map $\Phi$. Then $M^{\prime}$ is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, and the $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$-module
$\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}M^{\prime}$
is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}})^{K_{F^{\prime}}}$. We also have an isomorphism $\varphi:V_{F}^{K_{F}}\rightarrow V_{F^{\prime}}^{K_{F^{\prime}}}$, which is equivariant under the action of $\mathcal{H}(G_{F},K_{F})$ and hence $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$ via the Kazhdan isomorphism ${\rm Kaz}_{m}^{F}$. Using the isomorphisms $\Phi$, $\varphi$ and the relation (6.4), it follows from (6.7) that
$\mathcal{H}(G_{F,},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}M^{\prime}\simeq(V_{F^{\prime}}^{(l)})^{K_{F^{\prime}}}.$
This shows that the representation $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$. A similar argument also establishes the other implication, namely that the representation $\rho_{F}$ is linked with $\rho_{E}$ if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$. This completes the proof. ∎
Acknowledgements. The author is grateful to Santosh Nadimpalli for communicating Proposition 6.2, and for his helpful comments during the work.

### v4

@@
 Abstract.
-Let $G$ be a split connected reductive group, defined over $\mathbb{Z}$.
-Let $F$ and $F^{\prime}$ be the non-archimedean $m$-close local fields, where $m$ is a positive integer.
+Let $G$ be a connected split reductive group defined over $\mathbb{Z}$.
+Let $F$ and $F^{\prime}$ be two non-Archimedean $m$-close local fields, where $m$ is a positive integer.
 D.Kazhdan gave an isomorphism between the Hecke algebras ${\rm Kaz}_{m}^{F}:\mathcal{H}\big{(}G(F),K_{F}\big{)}\rightarrow\mathcal{H}\big{(}G(F^{\prime}),K_{F^{\prime}}\big{)}$, where $K_{F}$ and $K_{F^{\prime}}$ are the $m$-th usual congruence subgroups of $G(F)$ and $G(F^{\prime})$ respectively.
 On the other hand, if $\sigma$ is an automorphism of $G$ of prime order $l$, then we have Brauer homomorphism ${\rm Br}:\mathcal{H}(G(F),U(F))\rightarrow\mathcal{H}(G^{\sigma}(F),U^{\sigma}(F))$, where $U(F)$ and $U^{\sigma}(F)$ are compact open subgroups of $G(F)$ and $G^{\sigma}(F)$ respectively.
-This article aims to study the compatibility between these two maps in certain sense.
+In this article, we study the compatibility between these two maps in the local base change setting.
 Further, an application of this compatibility is given in the context of linkage–which is the representation theoretic version of Brauer homomorphism.
 1. Introduction
-The local Langlands correspondence relates the set of irreducible smooth representations of the group of rational points of a reductive group over a local field–with the representation theory of its Weil–Deligne group.
-For a split reductive group G defined over $\mathbb{Z}$, Kazhdan conjectured a link between the local Langlands correspondences for $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$, where $F$ is a $p$-adic field and $F^{\prime}$ is a non-Archimedean equal characteristic local field which are sufficiently close.
-Kazhdan’s approach is via isomorphisms between the respective Hecke algebras, with integral coefficients, $\mathcal{H}(\operatorname{{\bf G}}(F),K)$ and $\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K^{\prime})$, for some specific choice of open-compact subgroups $K$ and $K^{\prime}$ of $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$ respectively (see [Kaz86, Theorem A]).
+The local Langlands correspondence relates the set of irreducible smooth complex representations of the group of rational points of a reductive group over a local field–with the representation theory of its Weil–Deligne group.
+For a connected split reductive group G defined over $\mathbb{Z}$, D.
+Kazhdan conjectured a link between the local Langlands correspondences for $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$, where $F$ is a $p$-adic field and $F^{\prime}$ is a non-Archimedean positive characteristic local field–which are sufficiently close.
+Kazhdan’s approach is via isomorphisms between the respective Hecke algebras with complex coefficients, $\mathcal{H}(\operatorname{{\bf G}}(F),K)$ and $\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K^{\prime})$, for some specific choice of open-compact subgroups $K$ and $K^{\prime}$ of $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$ respectively (see [Kaz86, Theorem A]).
 Say $L$ is a non-Archimedean local field with residue characteristic $p$ and assume that $K$ is a nice compact open subgroup of $\operatorname{{\bf G}}(L)$–for instance a congruence subgroup of positive level.
-When $\operatorname{{\bf G}}(L)$ has an automorphism, denoted by $\sigma$, of prime order $l\neq p$ such that $\sigma(K)=K$, Treumann–Venkatesh ([TV16, Section 4]) define an $\overline{\mathbb{F}_{l}}$-algebra homomorphism between $\mathcal{H}(\operatorname{{\bf G}}(L),K)\otimes\overline{\mathbb{F}_{l}}$ and $\mathcal{H}(\operatorname{{\bf G}}^{\sigma}(L),K^{\sigma})\otimes\overline{\mathbb{F}_{l}}$, which is called the Brauer homomorphism.
-The Brauer homomorphism is conjectured to be compatible with Langlands functoriality principle.
-In this article, we study the compatibility of Brauer homomorhism and Kazhdan homomorphism in the Base change setting.
-Let $\mathfrak{o}_{L}$ be the ring of integers of $L$ and let $\mathfrak{p}_{L}$ be the maximal ideal of $\mathfrak{o}_{L}$.
-Let $E$ be a finite extension of $L$ of prime degree $l$, and let $\sigma$ be the generator of ${\rm Gal}(E/L)$.
-Let $K_{E}$ be the $m$-th congruence subgroup of $\operatorname{{\bf G}}(\mathfrak{o}_{E})$.
-Note that $K_{E}$ is stable under the action of $\sigma$ and let $K_{L}=K_{E}\cap\operatorname{{\bf G}}(L)$.
-Truemann and Venkatesh define Brauer homomorphism ${\rm Br}$, which is the algebra homomorphism
-${\rm Br}:\mathcal{H}(\operatorname{{\bf G}}(E),K_{E})^{\sigma}\rightarrow\mathcal{H}(\operatorname{{\bf G}}(L),K_{L})$
-obtained by restriction of functions on $\operatorname{{\bf G}}(E)$ to $\operatorname{{\bf G}}(L)$(see [TV16, Section 4]).
-The homomorphism ${\rm Br}$ is conjectured to be compatible with Langlands functoriality principle.
-Kazhdan ([Kaz86, Theorem A]) showed that there exists an integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close (see Section (4)), there is a Hecke algebra isomorphism
-${\rm Kaz}_{m}^{F}:\mathcal{H}(\operatorname{{\bf G}}(F),K_{F})\rightarrow\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K_{F^{\prime}}).$
-Kazhdan conjectured that it is sufficient to take $r=m$, and this conjecture is proved by Ganapathy in [Gan15].
-It is shown in [Gan15] that Kazhdan isomorphism is compatible with the various properties in representation theory, and it also establish a bridge between local Langlands correspondence(LLC) over the field characteristic zero and that of positive characteristic.
-We now set up some notations to formulate the main result of this article.
-Let $E$ be a finite Galois extension of a non-archimedean local field $F$ of prime degree $l$.
-Let $F^{\prime}$ be another non-Archimedean local field which is $m$-close to $F$.
-Throughout this article, we assume that the prime number $l$ is different from $p$, the residue characteristic of both $F$ and $F^{\prime}$.
-Lemma 4.2 in Section 4 says that we have a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ such that the fields $E$ and $E^{\prime}$ are $em$-close, where $e=1$ if $E$ is unramified and $e=l$ if $E$ is totally ramified.
+When $\operatorname{{\bf G}}(L)$ has an automorphism, denoted by $\sigma$, of prime order $l\neq p$ such that $\sigma(K)=K$, Treumann–Venkatesh ([TV16, Section 4]) define an $\overline{\mathbb{F}}_{l}$-algebra morphism between Hecke algebras with $\overline{\mathbb{F}}_{l}$-coefficients, $\mathcal{H}(\operatorname{{\bf G}}(L),K)^{\sigma}$ and $\mathcal{H}(\operatorname{{\bf G}}^{\sigma}(L),K^{\sigma})$, which is called the Brauer homomorphism, and it is conjectured to be compatible with Langlands functoriality.
+This article aims to study the compatibility between the Brauer homomorphism and Kazhdan isomorphism in the local base change setting.
+So far, this type of compatibility has not been addressed in the literature.
+Using this compatibility, one can relate the functoriality principles for connected reductive algebraic groups defined over the close local fields $F$ and $F^{\prime}$.
+This article gives one such treatment in the context of linkage (see [TV16, Section 6])–which is related to the functoriality over $\overline{\mathbb{F}}_{l}$.
+Let $\mathfrak{o}_{F}$ be the ring of integers of $F$, and let $\mathfrak{p}_{F}$ be the maximal ideal of $\mathfrak{o}_{F}$.
+Let $E$ be a Galois extension of $F$ of prime degree $l$, where $l$ is different from the residue characteristic of $F$.
+Let $\sigma$ be a generator of ${\rm Gal}(E/F)$.
+Similar notations follow for $E$.
+Let $K_{E}$ (resp.
+$K_{F}$) be the $m$-th congruence subgroup of $\operatorname{{\bf G}}(\mathfrak{o}_{E})$ (resp.
+$\operatorname{{\bf G}}(\mathfrak{o}_{F})$).
+Note that $K_{E}$ is stable under the action of $\sigma$ and let $K_{F}=K_{E}\cap\operatorname{{\bf G}}(F)$.
+Treumann–Venkatesh ( [TV16, Section 4]) define Brauer homomorphism, denoted by ${\rm Br}$–which is the $\overline{\mathbb{F}}_{l}$-algebra homomorphism
+${\rm Br}:\mathcal{H}(\operatorname{{\bf G}}(E),K_{E})^{\sigma}\rightarrow\mathcal{H}(\operatorname{{\bf G}}(F),K_{F}),$
+obtained by restriction of functions on $\operatorname{{\bf G}}(E)$ to $\operatorname{{\bf G}}(F)$.
+On the other hand, following Kazhdan’s approach in [Kaz86, Theorem A], we get an $\overline{\mathbb{F}}_{l}$-algebra isomorphism (Proposition 4.5)
+${\rm Kaz}_{m}^{F}:\mathcal{H}(\operatorname{{\bf G}}(F),K_{F})\rightarrow\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K_{F^{\prime}}),$
+provided the fields $F$ and $F^{\prime}$ are sufficiently close.
+The above isomorphism ${\rm Kaz}_{m}^{F}$ was originally constructed for complex Hecke algebras (Proposition 4.1)–which relies on the finiteness of complex Hecke algebras, due to Bernstein ([Ber84, Corollary 3.4]).
+For arbitrary Noetherian $\mathbb{Z}_{l}$-algebra $R$, where $\mathbb{Z}_{l}$ is the ring of integers of the field of $l$-adic numbers, the finiteness of Hecke algebras with coefficients in $R$ follows from the work of [DHKM24, Theorem 1.1].
+It enables us to define the isomorphism ${\rm Kaz}_{m}^{F}$ between mod-$l$ Hecke algebras.
+The method of close local fields is fruitful in obtaining analogous results for reductive groups defined over local fields of positive characteristic–which are known for reductive groups over $p$-adic fields.
+Using the close local fields approach, A.I.
+Badulescu ([Bad02]) proved the Jacquet-Langlands correspondence for the general linear group ${\rm GL}_{n}(F^{\prime})$.
+A similar technique was used by Badulescu–Henniart–Lemaire–Sécherre ([BHLS10]) to classify the smooth unitary dual of ${\rm GL}_{n}(D)$, where $D$ is a central division algebra over $F^{\prime}$.
+Kazhdan isomorphism and its various properties have been further studied by several others, notably by R.
+Ganapathy (see [Gan15], [Gan22]).
+We now state the main results of this article.
+Let G be a connected split reductive group defined over $\mathbb{Z}$.
+Let $F$ and $F^{\prime}$ be two non-Archimedean local fields with the same residue characteristic $p$.
+We assume that $F$ and $F^{\prime}$ are $m$-close (i.e., there exists a ring isomorphism $\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\xrightarrow{\sim}\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$).
+Let $E/F$ be a finite Galois extension of prime degree $l$ with $l\neq p$.
+Then there exists a finite Galois extension $E^{\prime}/F^{\prime}$ of degree $l$ such that $E^{\prime}$ is $em$-close to $E$, where $e=1$ if $E/F$ is unramified, and $e=l$ if $E/F$ is totally ramified (Lemma 4.6).
 Let $\sigma$ (resp.
-$\sigma^{\prime}$) be a generator of the cyclic group ${\rm Gal}(E/F)$ (resp.
+$\sigma^{\prime}$) be a generator of the group ${\rm Gal}(E/F)$ (resp.
 ${\rm Gal}(E^{\prime}/F^{\prime})$).
-This leads to the following diagram:
+Then we have the Kazhdan isomorphism ${\rm Kaz}_{em}^{E}$ between the mod-$l$ Hecke algebras $\mathcal{H}(\textbf{G}(E),K_{E})$ and $\mathcal{H}(\textbf{G}(E^{\prime}),K_{E^{\prime}})$, and it is equivariant under the Galois actions (Proposition 4.9)–which leads to the following diagram:
 […]
@@
 Theorem 1.1.
-Let $F$ and $F^{\prime}$ be the non-archimedean $m$-close local fields.
-Let $E$ be the finite Galois extensions of $F$ of prime degree $l$, and let $E^{\prime}$ be the Galois extension of $F^{\prime}$, as indicated above.
-Then for any split connected reductive group G defined over $\mathbb{Z}$, we have
+Let $F$ and $F^{\prime}$ be two non-Archimedean $m$-close local fields with residue characteristic $p$.
+Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$, and let $E^{\prime}$ be the Galois extension of $F^{\prime}$, as indicated above.
+Then, for any connected split reductive group G defined over $\mathbb{Z}$, we have
 ${\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E}={\rm Kaz}_{m}^{F}\circ{\rm Br}.$
-In the context of representation theory, there is a notion of linkage (see subsection 6.2, for definition), defined by Treumann-Venkatesh–which is the representation theoretic analouge of Brauer homomorphism in the sense that they are compatible ([TV16, Section 6]).
-In this article, we give a precise formulation (subsection 6.3) of the compatibility of Brauer homomorphism with linkage.
-Then, we use this formulation and Theorem 1.1 to establish the compatibility of linkage with Kazhdan isomorphism (Theorem 6.3).
-Now we briefly explain the contents of this article.
-In Section $2$, we review the Hecke algebra structure associated to any locally profinite group.
-In Section $3$, we formulate the Brauer homomorphism in our setting.
-The Kazdhan isomorphism and its compatibility with Galois actions will be discussed in Section $4$.
-In Section $5$, we prove Theorem 1.1.
-In Section $6$, we recall the notion of linkage and formulate some finiteness results on Tate cohomology group.
-Following this, we discuss the compatibility of linkage with Kazhdan isomorphism.
+Using the above theorem, we prove the compatibility of linkage with Kazhdan isomorphism.
+Say $(\pi,V)$ is an irreducible smooth $\overline{\mathbb{F}}_{l}$-representation of $\textbf{G}(E)$ such that $\pi$ is isomorphic to twisted representation $\pi^{\sigma}$.
+Then we have the Tate cohomology groups $\widehat{H}^{i}(\sigma,V)$, for $i\in\{0,1\}$, defined as $\overline{\mathbb{F}}_{l}$-representations of $\textbf{G}(F)$.
+An irreducible sub–quotient $\rho$ of $\widehat{H}^{i}(\sigma,V)$ is said to be linked with $\pi$.
+Let $\rho^{\prime}$ (resp.
+$\pi^{\prime}$) be the smooth irreducible $\overline{\mathbb{F}}_{l}$-representation of $\textbf{G}(F^{\prime})$ (resp.
+$\textbf{G}(E^{\prime})$) associated with $\rho$ (resp.
+$\pi$) via Kazhdan isomorphism and the natural correspondence between the set of simple modules over Hecke algebras and the set of irreducible representations of locally profinite group (see [Gan15, Section 2.3]).
+Then, using Theorem 1.1 and the compatibility of Brauer homomorphism with linkage (see subsection 6.4), we show that $\rho^{\prime}$ is linked with $\pi^{\prime}$ if and only if $\rho$ is linked with $\pi$ (Theorem 6.5).
+We now briefly explain the contents of this article.
+Section $2$ reviews the Hecke algebra structure associated with any locally profinite group.
+In Section $3$, we discuss the Brauer homomorphism.
+In Section $4$, we review the Kazdhan isomorphism and formulate its mod-$l$ version.
+We also set up some initial results, which are crucial for the main result.
+Theorem 1.1 is proved in Section $5$.
+In Section $6$, we prove the compatibility of linkage with the Kazhdan isomorphism.
+It also includes some finiteness results on the Tate cohomology groups.
 2. Hecke Algebra
-In this section, we review the Hecke algebra structure on any locally profinite group. We refer to [TV16, Subsection 2.10] for more insight.
+This section reviews the Hecke algebra structure on any locally profinite group. We refer to [TV16, Subsection 2.10] for more insight.
 2.1.
@@
 We denote by $G/K$ the set of all distinct left cosets of $K$ in $G$, and it is a discrete set with a left action of $G$.
-For any $g\in G$, the left coset $gK$ is sometimes denoted by the symbol $[g]$.
-There is a natural action of $G\times G$, via left translation, on the space of all $k$-valued functions on the discrete set $G/K\times G/K$.
+For any $g\in G$, we sometimes use the notation $[g]$ to denote the left coset $gK$.
+There is a natural action of $G\times G$, via left translation, on the space of $k$-valued functions on the discrete set $G/K\times G/K$.
 We denote by $\mathcal{F}(G//K)$ the space of all such $k$-valued functions on $G/K\times G/K$ that are invariant under the action of the diagonal subgroup $\Delta G=\{(g,g):g\in G\}$, and whose support is a union of finitely many $G$-orbits.
-There is a multiplicative structure on the space $\mathcal{F}(G//K)$, given by
+There is a multiplicative structure in the space $\mathcal{F}(G//K)$, given by
 $(f_{1}*f_{2})([x],[z])=\sum_{y\in G/K}f_{1}([x],[y])f_{2}([y],[z]),$
@@
 The space $\mathcal{F}(G//K)$ with the above multiplicative structure is a $k$-algebra, and it is called the Hecke algebra of $G$ associated with $K$.
-We now give an equivalent description of the Hecke algebra, which is relevant for our context.
+We now give an equivalent description of the Hecke algebra relevant to our context.
 2.2.
@@
 Let $\mathcal{H}(G,K)$ be the space of all $k$-valued functions which are bi-$K$-invariant i.e., $f(k_{1}gk_{2})=f(g),$ for all $k_{1},k_{2}\in K$, $g\in G$ and $f\in\mathcal{H}(G,K)$.
-Then the preceding bijection gives the following map
+Then, the preceding bijection gives the following map
 $\mathcal{F}(G//K)\longrightarrow\mathcal{H}(G,K)$
@@
 where the function $\widetilde{\phi}:G\rightarrow k$ is defined as $\widetilde{\phi}(g):=\phi([1],[g])$, for all $g\in G$.
-The map (2.1) is clearly a bijection, and this induces a $k$-algebra structure on the space $\mathcal{H}(G,K)$.
+The map (2.1) is a bijection, and it induces a $k$-algebra structure on the space $\mathcal{H}(G,K)$.
 In the literature, many authors refer to the $k$-algebra $\mathcal{H}(G,K)$ as the Hecke algebra of $G$ associated with $K$.
@@
     $gK^{\sigma}\mapsto gK$
-    is a bijection.
+    induces a bijection.
   2. (2)
-    There exists a subgroup $K^{\prime}$ of $K$ of finite index such that $K^{\prime}$ is an inverse limit of finite subgroups each of which is coprime to $l$.
+    There exists a subgroup $K^{\prime}$ of $K$ of finite index such that $K^{\prime}$ is an inverse limit of finite subgroups, each of which is coprime to $l$.
 3.2.
@@
 for all $x,y\in G$ and $\phi\in\mathcal{F}(G//K)$.
-It is shown in [TV16, Section 4] that the Brauer homomorphism ${\rm Br}$ is a $k$-algebra homomorphism from $\mathcal{F}(G//K)^{\sigma}$ to $\mathcal{F}(G^{\sigma}//K^{\sigma})$, and it is the restriction map
+In the article [TV16, Section 4], the authors proved that the Brauer homomorphism ${\rm Br}$ is a $k$-algebra homomorphism from $\mathcal{F}(G//K)^{\sigma}$ to $\mathcal{F}(G^{\sigma}//K^{\sigma})$, and defined by the restriction
 ${\rm Br}(\phi):={\rm Res}_{(G^{\sigma}/K^{\sigma}\times G^{\sigma}/K^{\sigma})}(\phi)$ (3.1)
@@
 $\sigma.(KgK):=K\sigma(g)K,$
-for all $g\in G$. Then we have a natural induced action of $\langle\sigma\rangle$ on the space $\mathcal{H}(G,K)$, defined as
+for all $g\in G$. Then, we have a natural induced action of $\langle\sigma\rangle$ on the space $\mathcal{H}(G,K)$, defined as
 $(\sigma.f)(g):=f(\sigma^{-1}(g)),$
-for all $g\in G$. Note that the bijection (2.1) preserves the action of the group $\langle\sigma\rangle$. Then we have commutative diagram
+for all $g\in G$. Note that the bijection (2.1) preserves the action of the group $\langle\sigma\rangle$. Then, we have a commutative diagram
 ${\mathcal{F}(G//K)^{\sigma}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm Br}}$${\mathcal{H}(G,K)^{\sigma}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm\overline{Br}}}$${\mathcal{F}(G^{\sigma}//K^{\sigma})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{H}(G^{\sigma},K^{\sigma})}$
-where the horizontal arrows in the diagram are the bijections induced by (2.1).
-We observe that the $k$-algebra map ${\rm\overline{Br}}:\mathcal{H}(G,K)^{\sigma}\rightarrow\mathcal{H}(G^{\sigma},K^{\sigma})$, induced by the above diagram, is the restriction map to subgroup $G^{\sigma}$ i.e.,
+Here, the horizontal arrows in the diagram are the bijections induced by (2.1).
+We observe that the $k$-algebra map ${\rm\overline{Br}}:\mathcal{H}(G,K)^{\sigma}\rightarrow\mathcal{H}(G^{\sigma},K^{\sigma})$, induced by the above diagram, is the restriction map to subgroup $G^{\sigma}$, i.e.,
 ${\rm\overline{Br}}(f):={\rm Res}_{G^{\sigma}}(f),$
@@
 4. Kazhdan isomorphism
-In this section, we review Kazhdan isomorphism and set up some initial results.
+In this part, we review Kazhdan isomorphism and set up some initial results.
 Let $F$ and $F^{\prime}$ be the non-Archimedean local fields, both of residue characteristic $p$, with ring of integers $\mathfrak{o}_{F}$ and $\mathfrak{o}_{F^{\prime}}$ and its maximal ideals $\mathfrak{p}_{F}$ and $\mathfrak{p}_{F^{\prime}}$ respectively.
@@
 Let $\textbf{B}=\textbf{T}\textbf{U}$ be a Levi decomposition over $\mathbb{Z}$ with maximal torus T and unipotent radical U.
-We denote by $X^{*}(\textbf{T})$ and $X_{*}(\textbf{T})$ the character lattice and the cocharacter lattice of T respectively.
-Then there is a natural $\mathbb{Z}$-bilinear pairing
+We denote by $X^{*}(\textbf{T})$ and $X_{*}(\textbf{T})$ the character lattice and the co-character lattice of T respectively.
+Then, there is a natural $\mathbb{Z}$-bilinear pairing
 $X^{*}(\textbf{T})\times X_{*}(\textbf{T})\longrightarrow\mathbb{Z}$
@@
 Let $\Phi\subseteq X^{*}(\textbf{T})$ be the set of roots of T in G, and let $\Phi^{+}$ be the set of positive roots of T in B.
-Let $G_{F}=\textbf{G}(F)$.
-Consider the subgroup
+4.1.1.
+Let $G_{F}$ denotes the group $\textbf{G}(F)$. Consider the subgroup
 $K_{F}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{F})\longrightarrow\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\big{)}.$
@@
 Fix a Haar measure $\mu_{F}$ on $G_{F}$ such that $\mu_{F}(K_{F})=1$.
-For $g\in G_{F}$, let $t_{g}=1_{K_{F}gK_{F}}$, the characteristic function on the double coset $K_{F}gK_{F}$.
+For each $g\in G_{F}$, we denote by $t_{g}$ the characteristic function on the double coset $K_{F}gK_{F}$.
 Then the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ is generated as a $k$-vector space by $\{t_{g}:g\in G_{F}\}$.
@@
 for all $x,y\in\textbf{G}(\mathfrak{o}_{F})$ and $g\in G_{F}(\mu)$.
-Then the finite group $H_{m}:=\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\times\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})$ acts transitively on the set $\{K_{F}xK_{F}:x\in G_{F}(\mu)\}$.
+Then, the finite group $H_{m}:=\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\times\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})$ acts transitively on the set $\{K_{F}xK_{F}:x\in G_{F}(\mu)\}$.
 Let $\Gamma_{\mu}(F)$ be the stabilizer of $K_{F}\varpi_{\mu}K_{F}$ in $H_{m}$.
@@
 $G_{F}=\coprod_{\mu\in X_{*}(\textbf{T})^{-}}\coprod_{(b_{i},b_{j})\in T_{\mu}(F)}K_{F}b_{i}\varpi_{\mu}b_{j}^{-1}K_{F}.$ (4.2)
-For each $\mu\in X_{*}(\textbf{T})^{-}$, the map $\Lambda$ induces a bijection $T_{\mu}(F)\rightarrow T_{\mu}(F^{\prime})$.
-Under this bijection, there is an isomorphism ([Kaz86, Section 1]) of $\mathbb{Z}$-modules:
+4.1.2.
+For each $\mu\in X_{*}(\textbf{T})^{-}$, the isomorphism $\Lambda$ induces a bijection $T_{\mu}(F)\rightarrow T_{\mu}(F^{\prime})$.
+Following Kazhdan’s constructions in [Kaz86, Section 1], we get an isomorphism of $\mathbb{Z}$-modules:
+$\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})\longrightarrow\mathcal{H}_{\mathbb{Z}}(G_{F^{\prime}},K_{F^{\prime}})$ (4.3)
+$t_{a_{i}\varpi_{\lambda}a_{j}^{-1}}\longmapsto t_{a_{i}^{\prime}\varpi_{\lambda}^{\prime}a_{j}^{\prime-1}},$
+and this induces a $\mathbb{C}$-vector space isomorphism
+${\rm Kaz}_{m,\mathbb{C}}^{F}:\mathcal{H}_{\mathbb{C}}(G_{F},K_{F})\longrightarrow\mathcal{H}_{\mathbb{C}}(G_{F^{\prime}},K_{F^{\prime}}),$
+where $\lambda\in X_{*}(\textbf{T})^{-}$, $(a_{i},a_{j})\in T_{\lambda}(F)$ and $(a_{i}^{\prime},a_{j}^{\prime})$ is the corresponding element in $T_{\lambda}(F^{\prime})$.
+The following result is due to Kazhdan ([Kaz86, Theorem A]).
+Proposition 4.1.
+Given $m\geq 1$.
+There exists a positive integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close, the map ${\rm Kaz}_{m,\mathbb{C}}^{F}$, defined above, is infact an isomorphism of $\mathbb{C}$-algebras.
+In order to establish that the vector space isomorphism ${\rm Kaz}_{m,\mathbb{C}}^{F}$ is compatible with algebra structures, the fields need to be a few levels closer.
+Kazhdan conjectured that one could take $r=m$ in Proposition 4.1–which was proved by R.
+Ganapathy ([Gan15, Corollary 3.15]).
+4.2. Kazhdan isomorphism over positive characteristic
+In this subsection, we formulate a version of Proposition 4.1 for mod-$l$ Hecke algebras.
+Recall that $\mathcal{H}(G_{F},K_{F})$ denotes the Hecke algebra, consisting of all $k$-valued smooth, compactly supported and bi-$K_{F}$-invariant functions on $G_{F}$.
+Suppose $F$ and $F^{\prime}$ are $m$-close.
+We denote by $X_{m}(F)$ the set of all $K_{F}$ double cosets in $G_{F}$.
+Similar notations are followed for $F^{\prime}$.
+The bijection $\phi:X_{m}(F)\rightarrow X_{m}(F^{\prime})$ gives the $\mathbb{Z}$-module isomorphism (see [Kaz86, Section 1]) of integral Hecke algebras
 $\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})\longrightarrow\mathcal{H}_{\mathbb{Z}}(G_{F^{\prime}},K_{F^{\prime}})$
-$t_{a_{i}\varpi_{\lambda}a_{j}^{-1}}\longmapsto t_{a_{i}^{\prime}\varpi_{\lambda}^{\prime}a_{j}^{\prime-1}},$
-and this induces a vector space isomorphism
-${\rm Kaz}_{m}^{F}:\mathcal{H}(G_{F},K_{F})\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}}),$
-where $\lambda\in X_{*}(\textbf{T})^{-}$, $(a_{i},a_{j})\in T_{\lambda}(F)$ and $(a_{i}^{\prime},a_{j}^{\prime})$ is the corresponding element in $T_{\lambda}(F^{\prime})$.
-For $F$ and $F^{\prime}$, sufficiently close fields, the following result is due to Kazhdan ([Kaz86, Theorem A]).
-Proposition 4.1.
-Given $m\geq 1$. There exists a positive integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close, the map ${\rm Kaz}_{m}^{F}$ defined above is an algebra isomorphism.
-Kazhdan conjectured that one could take $r=m$ in Proposition 4.1, as later proved by R.Ganapathy in her paper [Gan15].
-4.2.
-Let $F$ and $F^{\prime}$ be the non-Archimedean $m$-close local fields with residue chracteristic $p$. In [Del84], Deligne proved that there is an isomorphism
+$t_{x}\longmapsto t_{\phi(x)},$
+which induces an isomorphism of $k$-vector spaces
+${\rm Kaz}_{m}^{F}:\mathcal{H}(G_{F},K_{F})\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}}).$
+Kazhdan’s approach in proving the Hecke algebra isomorphism over $\mathbb{C}$, relies on the fact that $\mathcal{H}_{\mathbb{C}}(G_{F},K_{F})$ is finitely presented–which follows from the work of [Ber84].
+So, in order to establish that ${\rm Kaz}_{m}^{F}$ is an algebra map, we need the finiteness of Hecke algebras over $k$–which is made available due to seminal work of Dat–Helm–Kurinczuk–Moss ([DHKM24]).
+Lemma 4.2.
+For any compact open subgroup $K$ of $G_{F}$, the Hecke algebra $\mathcal{H}(G_{F},K)$ is finitely presented.
+Proof.
+Note that an algebra $\mathcal{A}$ is finitely presented if $\mathcal{A}$ is finitely generated and Noetherian. Then the above lemma follows from [DHKM24, Theorem 1.1 and Corollary 1.4]. ∎
+Remark 4.3.
+We now recall some results from [Kaz86]. Let $C$ be a finite subset of $X_{*}(\textbf{T})^{-}$ and let
+$G_{C}=\bigcup_{\mu\in C}G_{F}(\mu).$
+By [Kaz86, Lemma 1.3], there exists a positive integer $n_{C}\geq m$, such that for all $g\in G_{C}$, we have
+$gK_{n_{C}}(F)g^{-1}\subseteq K_{F},$
+where $K_{n_{C}}(F)$ is the $n_{C}$-th usual congruence subgroup of $G_{F}$. Moreover, if $F$ and $F^{\prime}$ are $n_{C}$-close, then
+${\rm Kaz}_{m}^{F}(f_{1}*f_{2})={\rm Kaz}_{m}^{F}(f_{1})*{\rm Kaz}_{m}^{F}(f_{2}),$ (4.4)
+for all $f_{1},f_{2}\in\mathcal{H}(G_{F},K_{F})$ supported on $G_{C}$.
+The following lemma, which was essential for [Kaz86, Theorem A], is valid over $k$ as well.
+Lemma 4.4.
+Let $C$ be a finite subset of $X_{*}(\textbf{T})^{-}$. Then
+  1. (1)
+    for any $\lambda,\mu\in X_{*}(\textbf{T})^{-}$, we have
+    $t_{\lambda(\varpi)}*t_{\mu(\varpi)}=t_{(\lambda+\mu)(\varpi)}$
+  2. (2)
+    for all $x_{1},x_{2}\in G_{C}$, we have
+    $t_{x_{1}\lambda(\varpi)x_{2}}=t_{x_{1}}*t_{\lambda(\varpi)}*t_{x_{2}}.$
+Proof.
+It follows from [All72, Theorem 2] by extension of scalars. ∎
+From these, we obtain Kazhdan isomorphism of mod-$l$ Hecke algebras. The proof exactly follows the same line of arguments of [Kaz86, Theorem A].
+Theorem 4.5.
+Given an integer $m\geq 1$. There exists a positive integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close, the map ${\rm Kaz}_{m}^{F}$ is an isomorphism of $k$-algebras.
+Proof.
+We will show that ${\rm Kaz}_{m}^{F}$ preserves the ring structures.
+Fix a finite subset $C\subseteq X_{*}(\textbf{T})^{-}$ containing 00, such that $C$ generates $X_{*}(\textbf{T})^{-}$ as a semigroup.
+Consider the finite set
+$X_{0}=\bigcup_{\lambda\in C}X_{\lambda},$
+where $X_{\lambda}=\big{\{}K_{F}gK_{F}\mid g\in\textbf{G}(\mathfrak{o}_{F})\lambda(\varpi)\textbf{G}(\mathfrak{o}_{F})\big{\}}$.
+It follows from Lemma 4.4 that the algebra $\mathcal{H}(G_{F},K_{F})$ is generated by the set $\big{\{}t_{g}:g\in X_{0}\big{\}}$, where $t_{g}$ denotes the characteristic function on the double coset $K_{F}gK_{F}$.
+Assume that $X_{0}=\{x_{1},\dots,x_{s}\}$.
+Let $k\langle X_{1},\dots,X_{s}\rangle$ denote the non-commutative polynomial algebra (free algebra) with generators $X_{1},\cdots,X_{s}$.
+Since the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ is finitely presented, there exists a surjective $k$-algebra homomorphism
+$\varphi:k\langle X_{1}\dots,X_{s}\rangle\rightarrow\mathcal{H}(G_{F},K_{F})$
+$X_{i}\longmapsto t_{x_{i}}$
+such that ${\rm Ker}(\varphi)$ is a two-sided ideal generated by non-commutative polynomials $f_{1},f_{2},\dots,f_{d}\in k\langle X_{1},\dots,X_{s}\rangle$. Let
+$M={\rm max}\big{\{}a_{i}:1\leq i\leq d\big{\}},$
+where $a_{i}$ is the degree of the polynomial $f_{i}$. Then there exists a finite set $B\subseteq\Phi^{+}$ such that
+$X_{B}=\bigcup_{\lambda\in B}X_{\lambda}$
+contains a compact set of the form $\big{\{}g_{1}g_{2}\dots g_{M}:g_{i}\in X_{0}\big{\}}$. By Remark 4.3, there is an integer $r=r_{B}\geq m$, such that if $F$ and $F^{\prime}$ are $r$-close, we have
+${\rm Kaz}_{m}^{F}(f_{1}*f_{2})={\rm Kaz}_{m}^{F}(f_{1})*{\rm Kaz}_{m}^{F}(f_{2}),$
+for all functions $f_{1},f_{2}$ supported on $G_{B}$. Now, for each $1\leq i\leq d$, the above identity gives
+$f_{i}\big{(}{\rm Kaz}_{m}^{F}(t_{x_{1}}),\dots,{\rm Kaz}_{m}^{F}(t_{x_{s}})\big{)}=0$
+or, equivalently
+$f_{i}\big{(}t_{\phi(x_{1})},\dots,t_{\phi(x_{s})}\big{)}=0.$
+Consider the homomorphism
+$\varphi^{\prime}:k\langle X_{1},\dots,X_{s}\rangle\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$
+$X_{j}\longmapsto t_{\phi(x_{j})}.$
+Note that $\varphi^{\prime}$ factors through ${\rm Ker}(\varphi)$, and hence we get a unique algebra morphism
+$\widetilde{\varphi^{\prime}}:\mathcal{H}(G_{F},K_{F})\longrightarrow\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$
+such that $\widetilde{\varphi^{\prime}}(t_{x_{j}})=t_{\phi(x_{j})}$, for each $j$. Since $C$ generates $X_{*}(\textbf{T})^{-}$ as a semigroup, using the identities in Lemma 4.4, we get that
+$\widetilde{\varphi^{\prime}}(t_{x})=t_{\phi(x)}={\rm Kaz}_{m}^{F}(t_{x}),$
+for all $x\in X_{m}(F)$. This shows that $\widetilde{\varphi^{\prime}}={\rm Kaz}_{m}^{F}$, which is an algebra morphism. Hence, the theorem. ∎
+4.3. Galois extension of close local fields
+Let $F$ and $F^{\prime}$ be two non-Archimedean $m$-close local fields with residue characteristic $p$. Deligne ([Del84, Section 3.5]) proved that there is an isomorphism
 ${\rm Gal}(\overline{F}/F)/\mathcal{I}_{F}^{m}\xrightarrow{{\rm Del}_{m}}{\rm Gal}(\overline{F^{\prime}}/F^{\prime})/\mathcal{I}_{F^{\prime}}^{m},$
@@
 Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$.
-Following the description of the map ${\rm Del}_{m}$, as mentioned in [Gan15, Section 2.1], we have a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of same degree $l$.
-Moreover, the extension $E^{\prime}/F^{\prime}$ can be constructed so that $E$ and $E^{\prime}$ are likewise sufficiently close, as ensured by the next lemma.
+Following the description of the isomorphism ${\rm Del}_{m}$ (see [Del84, Section 1.3] and [Gan15, Section 2.1]), we have a finite Galois extension $E^{\prime}/F^{\prime}$ of same degree $l$.
+Moreover, the extension $E^{\prime}/F^{\prime}$ can be so constructed that the fields $E$ and $E^{\prime}$ are also sufficiently close.
 Lemma 4.6.
-Let $F$ and $F^{\prime}$ be the $m$-close non-Archimedean local fields, and let $E$ be the finite Galois extension of $F$ of prime degree $l$.
-Then there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ such that $E$ and $E^{\prime}$ are $em$-close, where $e\in\{1,l\}$.
+[…]
 Proof.
-The proof is divided into two parts.
-In the first part, we consider the case when $E$ is unramified over $F$.
+The lemma follows from [Del84].
+For completeness, we give a sketch of the proof here.
+The proof is in two parts.
+In the first part, we consider the situation where $E$ is unramified over $F$.
 The second part deals with totally ramified case.
@@
 Let $E^{\prime}$ be the field extension $F^{\prime}[T]/(f^{\prime})$.
-Then $E^{\prime}$ is unramified over $F^{\prime}$ and there is a ring isomorphism
+Then $E^{\prime}$ is unramified over $F^{\prime}$, and there is a ring isomorphism
 $\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{m}\simeq A_{m}^{\prime}[T]/(\overline{f}^{\prime}).$
@@
 Recall that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the isomorphism $\Lambda$.
-Now consider the field extension
+Now consider the field extension,
 $E^{\prime}:=F[T]/(T^{l}-\varpi^{\prime})$
@@
 Let $\overline{\varpi}$ be the image of $\varpi$ under ${\rm mod}-\mathfrak{p}_{F}^{m}$.
-Then the natural surjection
+Then, the natural surjection
 $\mathfrak{o}_{F}[T]/(T^{l}-\varpi)\longrightarrow A_{m}[T]/(T^{l}-\overline{\varpi})$
-factorzies through the ideal $\mathfrak{p}_{F}^{m}[T]/(T^{l}-\varpi)$, and it gives the ring isomorphism
+factorizes through the ideal $\mathfrak{p}_{F}^{m}[T]/(T^{l}-\varpi)$, and it gives the ring isomorphism
 $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{lm}\xrightarrow{\sim}A_{m}[T]/(T^{l}-\overline{\varpi}).$
-Similarly, there is a ring isomorphism
+Similarly, there is a ring isomorphism,
 $\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{lm}\simeq A_{m}^{\prime}[T]/(T^{l}-\overline{\varpi^{\prime}}),$
@@
 ∎
-4.2.3.
-For each $e\in\{1,l\}$, we fix a ring isomorphism $\Pi:\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\rightarrow\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}$ as above, induced by $\Lambda$, so that the diagram
+Remark 4.7.
+Let $F$, $F^{\prime}$, $E$ and $E^{\prime}$ be as in Lemma 4.6.
+We simply write $E$ and $E^{\prime}$ are $em$-close, where we define $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified.
+We fix a ring isomorphism $\Pi:\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\rightarrow\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}$ as above, induced by $\Lambda$, which makes commutative the diagram
 ${\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Lambda}$${\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}_{E^{\prime}}^{em}}$
-is commutative, where the vertical arrows are the natural inclusion maps.
+Here, the vertical arrows are the natural inclusion maps.
 We choose uniformizer $\pi$ (resp.
@@
 For any $\overline{a}\in\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}$ (or, $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}$), we write $\overline{a}^{\prime}$ for the image $\Lambda(\overline{a})$ (or, $\Pi(\overline{a})$), whenever no confusion arises.
+Remark 4.8.
+Let $\Gamma$ and $\Gamma^{\prime}$ denote the Galois groups ${\rm Gal}(E/F)$ and ${\rm Gal}(E^{\prime}/F^{\prime})$ respectively.
+We choose generators $\sigma$ and $\sigma^{\prime}$ of $\Gamma$ and $\Gamma^{\prime}$ respectively, which makes commutative the following diagram
+${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$$\scriptstyle{\sigma}$${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma^{\prime}}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}}$
+We now describe the generators $\sigma$ and $\sigma^{\prime}$ in totally ramified case.
+For the unramified case, one similarly chooses $\sigma$ and $\sigma^{\prime}$.
+So, assume that $E/F$ is totally ramified.
+Then, Lemma 4.6 says that $E^{\prime}/F^{\prime}$ is also totally ramified and $E$ is $lm$-close to $E^{\prime}$.
+Since $l\neq p$, the extension $E/F$ is tamely ramified, and we have
+$E=\dfrac{F[T]}{(T^{l}-\varpi)}$ and $E^{\prime}=\dfrac{F^{\prime}[T]}{(T^{l}-\varpi^{\prime})}$.
+From (4.3.2) of Lemma 4.6, we get the following isomorphism
+$\Pi:\frac{A_{m}[T]}{(T^{l}-\overline{\varpi})}\longrightarrow\frac{A_{m}^{\prime}[T]}{(T^{l}-\overline{\varpi^{\prime}})},$
+induced by the ring isomorphism $\Lambda:A_{m}\rightarrow A_{m}^{\prime}$ and it is given by
+$\Pi\Big{(}\sum_{i=0}^{l-1}\overline{x_{i}}T^{i}+(T^{l}-\overline{\varpi})\Big{)}=\sum_{i=0}^{l-1}\Lambda(\overline{x_{i}})T^{i}+(T^{l}-\overline{\varpi^{\prime}}),$
+where $\overline{x_{i}}\in A_{m}$, for each $i$.
+Let $\sigma$ and $\sigma^{\prime}$ be the automorphisms of $E$ and $E^{\prime}$ respectively, that sends the class of $T$ to the class of $T^{2}$, i.e.,
+$\sigma(T+(T^{l}-\varpi))=T^{2}+(T^{l}-\varpi)$
+and
+$\sigma^{\prime}(T+(T^{l}-\varpi^{\prime}))=T^{2}+(T^{l}-\varpi^{\prime}),$
+Then $\sigma$ (resp.
+$\sigma^{\prime}$) is a generator of $\Gamma$ (resp.
+$\Gamma^{\prime}$).
+Moreover, the automorphism $\sigma$ induces a ring isomorphism $\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}\xrightarrow{\sim}\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em}$, and the induced map is again denoted by $\sigma$.
+Similarly for $\sigma^{\prime}$.
+Then we have
+$\Pi\circ\sigma=\sigma^{\prime}\circ\Pi.$
 4.4. Compatibility with Galois action
-Here, we continue to work with the set up of the preceeding subsection.
-We fix a generator $\sigma$ of the group ${\rm Gal}(E/F)$.
-Then $\sigma$ can be regarded as an automorphism of the group $G_{E}$, via its natural action on $G_{E}$.
-Consider the compact open subgroup
+Here, we continue to work with the setup of the preceding subsection.
+We choose the generator $\sigma$ (resp.
+$\sigma^{\prime}$) of $\Gamma$ (resp.
+$\Gamma^{\prime}$), as discussed in Remark 4.8.
+We have the following commutative diagram
+${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$$\scriptstyle{\sigma}$${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma^{\prime}}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}}$
+where $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified. To be more precise, for any $a\in\mathfrak{o}_{E}$, we have the relation
+$\overline{\sigma(a)}^{\prime}=\sigma(\overline{a})^{\prime}=\sigma^{\prime}(\overline{a}^{\prime})=\overline{\sigma^{\prime}(a^{\prime})},$ (4.5)
+where $a^{\prime}\in\mathfrak{o}_{E^{\prime}}$ is the preimage of $\overline{a}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$.
+We fix an isomorphism $\iota:\Gamma\xrightarrow{\sim}\Gamma^{\prime}$ that sends the generator $\sigma$ to $\sigma^{\prime}$.
+4.5.
+Note that $\sigma$ can be regarded as an automorphism of the group $G_{E}$ via its natural action on $G_{E}$. Consider the compact open subgroup,
 $K_{E}={\rm Ker}\big{(}\textbf{G}(\mathfrak{o}_{E})\rightarrow\textbf{G}(\mathfrak{o}_{E}/\mathfrak{p}_{E}^{em})\big{)}.$
-We observe that $K_{E}$ is $\sigma$-plain, and there is an induced action of $\langle\sigma\rangle$ on the Hecke algebra $\mathcal{H}(G_{E},K_{E})$, given by
+Since $K_{E}$ is a pro-$p$ subgroup of $G_{E}$, it follows from the arguments of [Fen24, Lemma 6.6] that $K_{E}$ is $\sigma$-plain.
+There is an induced action of $\langle\sigma\rangle$ on the Hecke algebra $\mathcal{H}(G_{E},K_{E})$, given by
 $(\sigma.f)(x)=f(\sigma^{-1}x),$
@@
 Similarly, there is an action of $\langle\sigma^{\prime}\rangle$ on the Hecke algebra $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$.
-We choose a group isomorphism $\iota:{\rm Gal}(E/F)\rightarrow{\rm Gal}(E^{\prime}/F^{\prime})$, sending $\sigma$ to a generator, say $\sigma^{\prime}$, that makes commutative the diagram
-${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$$\scriptstyle{\sigma}$${\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma^{\prime}}$${\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}}}$
-where the horizontal maps are the ring isomorphisms induced by $\sigma:E\rightarrow E$ and $\sigma^{\prime}:E^{\prime}\rightarrow E^{\prime}$.
-To be more precise, for any $a\in\mathfrak{o}_{E}$, we have the relation
-$\overline{\sigma(a)}^{\prime}=\sigma(\overline{a})^{\prime}=\sigma^{\prime}(\overline{a}^{\prime})=\overline{\sigma^{\prime}(a^{\prime})},$ (4.3)
-where $a^{\prime}\in\mathfrak{o}_{E^{\prime}}$ is the preimage of $\overline{a}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$.
 We end this section with the following proposition, which shows the compatibility of Kazhdan isomorphism with these Galois actions.
@@
 Proof.
-In view of the Cartan decomposition (4.2), it is enough to show that
+Because of the Cartan decomposition (4.2), it is enough to show that
 ${\rm Kaz}_{em}^{E}(\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}})=\sigma^{\prime}.{\rm Kaz}_{em}^{E}(t_{a_{i}\pi_{\mu}a_{j}^{-1}}),$
@@
 $\sigma.t_{a_{i}\pi_{\mu}a_{j}^{-1}}=t_{\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})}.$
-By Cartan decomposition (4.2), the double coset $K_{E}\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})K_{E}$ is of the form $K_{E}c_{i}\pi_{\mu}c_{j}^{-1}K_{E}$, where $(c_{i},c_{j})\in T_{\mu}(E)$.
+By Cartan decomposition (4.2), the double coset space $K_{E}\sigma(a_{i})\sigma(\pi_{\mu})\sigma(a_{j}^{-1})K_{E}$ is of the form $K_{E}c_{i}\pi_{\mu}c_{j}^{-1}K_{E}$, for some $(c_{i},c_{j})\in T_{\mu}(E)$.
 Then
@@
 Let $d_{j}^{\prime}\in\textbf{G}(\mathfrak{o}_{E^{\prime}})$ be a preimage of $\overline{d_{j}}^{\prime}$ under ${\rm mod}-\mathfrak{p}_{E^{\prime}}^{em}$.
-Then the commutative diagram
+Then, the commutative diagram
 ${\mathbb{G}_{m}(\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$$\scriptstyle{\mu}$${\textbf{T}(\mathfrak{o}_{E}/\mathfrak{p}^{em}_{E})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${\mathbb{G}_{m}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$${\textbf{T}(\mathfrak{o}_{E^{\prime}}/\mathfrak{p}^{em}_{E^{\prime}})}$
@@
 $K_{E^{\prime}}c_{i}^{\prime}\pi_{\mu}^{\prime}c_{j}^{\prime-1}K_{E^{\prime}}$ $=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\pi_{\mu}^{\prime}\sigma^{\prime}(d_{j}^{\prime-1})K_{E^{\prime}}$
-$=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})K_{E^{\prime}}.$
-This shows that
+$=K_{E^{\prime}}\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})K_{E^{\prime}},$
+which shows that
 $t_{c_{i}^{\prime}\pi_{\mu}^{\prime}c_{j}^{\prime-1}}=t_{\sigma^{\prime}(a_{i}^{\prime})\sigma^{\prime}(\pi_{\mu}^{\prime})\sigma^{\prime}(a_{j}^{\prime-1})}=\sigma^{\prime}.t_{a_{i}^{\prime}\pi_{\mu}^{\prime}a_{j}^{\prime-1}}.$
@@
 5. Compatibility of Brauer homomorphism and Kazhdan isomorphism
-In this section, we prove Theorem 1.1. Let us first recall the notations and terminologies from the prceeding sections.
+In this section, we prove Theorem 1.1. Let us first recall the notations and terminologies from the preceding sections.
 5.1.
@@
 Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$, where $p$ is the residue characteristic of both $F$ and $F^{\prime}$.
-According to Lemma 4.2, there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ and $E^{\prime}$ is $em$-close to $E$, where $e\in\{1,l\}$.
+According to Lemma 4.6, there exists a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ and $E^{\prime}$ is $em$-close to $E$, where $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified.
 Let G be a split connected reductive group defined over $\mathbb{Z}$.
@@
 Theorem 5.1.
-Let $F,F^{\prime},E$ and $E^{\prime}$ be as above. Then for any split connected reductive group G defined over $\mathbb{Z}$, we have
-${\rm Kaz}_{m}^{F}\circ\overline{\rm Br}=\overline{\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E}.$
+Let $F,F^{\prime},E$ and $E^{\prime}$ be as above. Then, for any connected split reductive group G defined over $\mathbb{Z}$, we have
+${\rm Kaz}_{m}^{F}\circ\overline{\rm Br}=\overline{\rm Br}^{\prime}\circ{\rm Kaz}_{em}^{E},$
+where $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified.
 Proof.
-The proof is divided into two parts.
-In the first part, we consider the extension $E/F$ to be unramified.
-The second part deals with totally ramified case.
-To start with, let us recall that an arbitrary $h\in\mathcal{H}(G_{E},K_{E})^{\sigma}$ can be written as
+We divide the proof into two parts.
+In the first part, we consider the extension $E/F$ unramified.
+The second part deals with the totally ramified case.
+To begin with, let us recall that an arbitrary $h\in\mathcal{H}(G_{E},K_{E})^{\sigma}$ can be written as
 $h=\sum_{\mu\in X_{*}(\textbf{T})^{-}}\sum_{(a_{i},a_{j})\in T_{\mu}(E)}\alpha_{\mu}(a_{i},a_{j})t_{a_{i}\pi_{\mu}a_{j}^{-1}},$
@@
 $\varpi_{\xi}\in K_{E}c_{i}^{-1}a_{i}\pi_{\mu}a_{j}^{-1}c_{j}K_{E},$
-and it is possible only if $\xi^{l}=\mu$. Therefore, it follows that
-$\pi_{\xi^{l}}\in K_{E}c_{i}^{-1}a_{i}\pi_{\xi^{l}}a_{j}^{-1}c_{j}K_{E}$
+and it is possible only if $l\xi=\mu$. Therefore, it follows that
+$\pi_{l\xi}\in K_{E}c_{i}^{-1}a_{i}\pi_{l\xi}a_{j}^{-1}c_{j}K_{E}$
 if and only if
-$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{\xi^{l}}(E).$
-Under the bijection $\Gamma_{\xi^{l}}(E)\rightarrow\Gamma_{\xi^{l}}(E^{\prime})$, we have
-$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{\xi^{l}}(E^{\prime})$
+$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{l\xi}(E).$
+Under the bijection $\Gamma_{l\xi}(E)\rightarrow\Gamma_{l\xi}(E^{\prime})$, we have
+$(\overline{c_{i}}^{-1}\overline{a_{i}},\overline{a_{j}}^{-1}\overline{c_{j}})\in\Gamma_{l\xi}(E^{\prime})$
 if and only if
-$(\overline{c_{i}}^{\prime-1}\overline{a_{i}}^{\prime},\overline{c_{j}}^{\prime-1}\overline{a_{j}}^{\prime})=(\overline{c_{i}^{\prime}}^{-1}\overline{a_{i}^{\prime}},\overline{c_{j}^{\prime}}^{-1}\overline{a_{j}^{\prime}})\in\Gamma_{\xi^{l}}(E^{\prime}),$
-where $(c_{i}^{\prime},c_{j}^{\prime})\in T_{\xi^{l}}(F^{\prime})$. Thus we have
+$(\overline{c_{i}}^{\prime-1}\overline{a_{i}}^{\prime},\overline{c_{j}}^{\prime-1}\overline{a_{j}}^{\prime})=(\overline{c_{i}^{\prime}}^{-1}\overline{a_{i}^{\prime}},\overline{c_{j}^{\prime}}^{-1}\overline{a_{j}^{\prime}})\in\Gamma_{l\xi}(E^{\prime}),$
+where $(c_{i}^{\prime},c_{j}^{\prime})\in T_{l\xi}(F^{\prime})$. Thus we have
 $c_{i}\varpi_{\xi}c_{j}^{-1}\in K_{E}a_{i}\pi_{\mu}a_{j}^{-1}K_{E}\iff c_{i}^{\prime}\varpi^{\prime}_{\xi}c_{j}^{\prime-1}\in K_{E^{\prime}}a_{i}^{\prime}\pi^{\prime}_{\mu}a_{j}^{\prime-1}K_{E^{\prime}}.$ (5.7)
@@
 6. Linkage over close local fields
-In this section, an application of Theorem 5.1 is given in the context of linkage.
-To be more precise, we show that the linkage is compatible with Kazhdan isomorphism.
-To prove this, we first give a formulation of the compatibility of linkage with Brauer homomorphism, as described in [TV16, Section 6.2], and then we use Theorem 5.1 to get the required compatibility.
-To start with, we first recall the Tate cohomology and the concept of linkage.
-For a reference, see [TV16, Section 3 and 6].
+This section gives an application of Theorem 5.1 in the context of representation theory.
+In the article [TV16], Treumann–Venkatesh defined the notion of linkage, which is conjectured to be compatible with the Langlands functorial transfer.
+Using Theorem 5.1, we show that linkage is compatible under close local fields.
+To prove this, we first formulate the compatibility of linkage with Brauer homomorphism, as described in [TV16, Section 6.2].
+To begin with, we recall the notion of linkage and Tate cohomology.
+For precise reference, see [TV16, Section 3 and 6].
 6.1. Tate cohomology
-Let $\Gamma$ be a cyclic group of order $n$ with generator $\gamma$. Let $M$ be an abelian group with an action of $\Gamma$. Let $T_{\gamma}$ be the automorphism of $M$ defined by
+Let $\Gamma$ be a cyclic group of order $n$ with a generator $\gamma$, and let $M$ be an abelian group with an action of $\Gamma$. Let $T_{\gamma}$ be the automorphism of $M$ defined by
 $T_{\gamma}(m)=\gamma.m$, for $\gamma\in\Gamma,m\in M$.
-Let $N={\rm id}+T_{\gamma}+T_{\gamma^{2}}+\dots+T_{\gamma^{n-1}}$ be the norm operator. The Tate cohomology groups $\widehat{H}^{0}(M)$ and $\widehat{H}^{1}(M)$ are defined as
+Let $N={\rm id}+T_{\gamma}+T_{\gamma^{2}}+\dots+T_{\gamma^{n-1}}$ be the norm operator. The Tate cohomology groups $\widehat{H}^{i}(M)$, for $i\in\{0,1\}$, are defined as
 $\widehat{H}^{0}(M):={\rm Ker}({\rm id}-T_{\gamma})/{\rm Im}(N)$,
 $\widehat{H}^{1}(M):={\rm Ker}(N)/{\rm Im}({\rm id}-T_{\gamma})$.
-6.1.1. Tate Cohomology of sheaves on $\ell$-spaces
-Let $X$ be an $\ell$-space (i.e., Hausdorff, locally compact and totally disconnected) with an action of a finite group $\langle\gamma\rangle$ of prime order $l$.
+6.1.1. Tate Cohomology of sheaves on $l$-spaces
+Let $X$ be an $l$-space (i.e., locally compact, totally disconnected and Hausdorff) with an action of a finite group $\langle\gamma\rangle$ of prime order $l$.
 Let $\mathcal{F}$ be a sheaf of $k$-modules on $X$.
@@
 $\widehat{H}^{1}(\mathcal{F}|_{X^{\gamma}}):={\rm Ker}(N)/{\rm Im}(1-\gamma)$.
-For each $i\in\{0,1\}$, the object $\widehat{H}^{i}(\mathcal{F}|_{X^{\gamma}})$ is a sheaf on $X^{\gamma}$.
+For each $i$, the object $\widehat{H}^{i}(\mathcal{F}|_{X^{\gamma}})$ is a sheaf on $X^{\gamma}$.
+If $\Gamma_{c}(X;\mathcal{F})$ denotes the space of compactly supported sections on $\mathcal{F}$, then we have the following result.
+Proposition 6.1.
+[TV16, Proposition 3.3.1] For each $i\in\{0,1\}$, the restriction map induces an isomorphism
+$\widehat{H}^{i}(\Gamma_{c}(X;\mathcal{F}))\xrightarrow{\sim}\Gamma_{c}(X^{\gamma};\widehat{H}^{i}(\mathcal{F}))$.
 6.2. Linkage
-Let $F$ be a non-archimedean local field and let $G$ be the $F$-points of a connected reductive group defined over $F$.
+Let $F$ be a non-Archimedean local field, and $G$ be the $F$-points of a connected reductive group defined over $F$.
 A representation $(\rho,V)$ of $G$ is called smooth if for every vector $v\in V$, the $G$-stabilizer of $v$ is an open subgroup of $G$.
-Moreover, such a representation $\rho$ is called admissible if for each compact open subgroup $K$ of $G$, the space of $K$-fixed vectors $V^{K}$ (sometimes denoted by $\rho^{K}$) is finite-dimensional over $k$.
-In this subsection, all the representations are assumed to be admissible and the representation spaces are defined over $k$.
-We denote by $\mathcal{R}(G)$ the category of smooth admissible $k$-representations of $G$.
+The representation $\rho$ is called admissible if, for each compact open subgroup $K$ of $G$, the space of $K$-fixed vectors $V^{K}$ (sometimes denoted by $\rho^{K}$) is finite-dimensional over $k$.
+In this subsection, all the representations are assumed to be admissible, and the representation spaces are defined over $k$.
+We denote by $\mathcal{R}(G)$ the category of smooth, admissible, $k$-representations of $G$.
 6.2.1. Frobenius Twist
@@
 $c\star v=c^{\frac{1}{l}}.v$ , for all $c\in k,v\in V$.
-Then the action of $G$ on $V$ induces a representation $\rho^{(l)}$ of $G$ on $V^{(l)}$. The representation $(\rho^{(l)},V^{(l)})$ is called the Frobenius twist of the representation $(\rho,V)$.
+Then, the action of $G$ on $V$ induces a smooth representation $\rho^{(l)}$ of $G$ on the space $V^{(l)}$. The representation $(\rho^{(l)},V^{(l)})$ is called the Frobenius twist of $(\rho,V)$.
 6.2.2.
 Let $\sigma$ be an automorphism of $G$ of prime order $l$.
-An irreducible representation $(\Xi,W)$ of $G$ is called $\sigma$-fixed if $\Xi$ is isomorphic to $\Xi\circ\sigma$.
-For such representation $\Xi$, there is an unique action of $\sigma$ on the space $W$, which is compatible with the action of $\sigma$ on $G$ ([TV16, Proposition 6.1]).
-Then we have the Tate cohomology space $\widehat{H}^{i}(W)$, defined as $k$-representation of $G^{\sigma}$.
-We denote this representation by $\widehat{H}^{i}(\Xi)$.
+An irreducible representation $(\Xi,W)$ of $G$ is called $\sigma$-fixed if $\Xi$ is isomorphic to the twisted representation $\Xi\circ\sigma$.
+In such a case, there is a unique action of $\sigma$ on $W$, compatible with the action of $\sigma$ on $G$ (see [TV16, Proposition 6.1]).
+Then the Tate cohomology groups $\widehat{H}^{i}(W)$, for $i\in\{0,1\}$, are defined as $k$-representations of $G^{\sigma}$.
+We denote these representations by $\widehat{H}^{i}(\Xi)$.
 Definition 6.2.
-An irreducible representation $\rho$ of $G^{\sigma}$ is linked with the representation $\Xi$ if the Frobenius-twist $\rho^{(l)}$ occurs as a Jordan-Holder constituent of $\widehat{H}^{0}(\Xi)$ or $\widehat{H}^{1}(\Xi)$.
-6.2.3.
-It was conjectured by Venkatesh-Treumann ([TV16, Section 6.3]) that $\widehat{H}^{i}(\Xi)$ is of finite length as a representation of $G^{\sigma}$.
-We give a proof of this conjecture for the group ${\rm GL}_{n}$, under the following set up.
-[…]
-We start with a non-archimedean local field $L$. Let $P_{n}(L)$ and $U_{n}(L)$ be the subgroups of ${\rm GL}_{n}(L)$, given by
+An irreducible representation $\rho$ of $G^{\sigma}$ is linked with the representation $\Xi$ if the Frobenius twist $\rho^{(l)}$ occurs as a Jordan-Holder constituent of $\widehat{H}^{0}(\Xi)$ or $\widehat{H}^{1}(\Xi)$.
+In [TV16, Section 6.3], the authors made a conjecture that the Tate cohomology groups $\widehat{H}^{i}(\Xi)$, $i\in\{0,1\}$, are of finite length as representations of $G^{\sigma}$.
+We give a proof of this conjecture for the general linear group ${\rm GL}_{n}$ in the local base change setting, using Bernstein–Zelevinksy derivatives.
+6.3. Finiteness of Tate cohomology
+Before going into the main theorem, we first recall the notion of derivatives for smooth representations of ${\rm GL}_{n}(L)$, where $L$ is a non-Archimedean local field.
+We refer to [Vig96, Chapter III] and [BZ77, Section 3] for more details.
+Let $P_{n}(L)$ and $U_{n}(L)$ be the subgroups of ${\rm GL}_{n}(L)$, given by
 $P_{n}(L)=\Big{\{}\begin{pmatrix}A&B\\ 0&1\end{pmatrix}:A\in{\rm GL}_{n-1}(L),B\in L^{n-1}\Big{\}}$
@@
 $U_{n}(L)=\Big{\{}\begin{pmatrix}1&C\\ 0&1\end{pmatrix}:C\in L^{n-1}\Big{\}}$
-respectively. We denote by $X_{L}$ the coset space $P_{n}(L)/U_{n}(L)$. Let $\theta_{n}$ be the character of $U_{n}(L)$, defined as
-$\theta_{n}((u)_{i,j}):=\psi_{L}(u_{n-1,n})$, for $(u)_{i,j}\in U_{n}(L)$.
-We have four fundamental functors:
+respectively.
+We use the short notation $X_{L,n}$ to denote the coset space $P_{n}(L)/P_{n-1}(L)U_{n}(L)$.
+Let $\psi_{L}:L\rightarrow k^{\times}$ be a non-trivial additive character.
+Let $N_{n}(L)$ be the group of all unipotent upper triangular matrices in ${\rm GL}_{n}(L)$.
+Let $\Theta_{L}$ be the character of $N_{n}(L)$, defined as
+$\Theta_{L}((a_{ij})_{i,j=1}^{n}):=\psi_{L}(a_{12}+a_{23}+\cdots a_{n-1,n}),$ (6.1)
+for $(a_{ij})_{i,j=1}^{n}\in N_{n}(L)$. By abuse of notation, the restriction ${\rm res}_{U_{n}(L)}\Theta_{L}$ is also denoted by $\Theta_{L}$. We have four fundamental functors:
 $\Psi^{-}:\mathcal{R}(P_{n}(L))\rightarrow\mathcal{R}({\rm GL}_{n-1}(L)),\Psi^{+}:\mathcal{R}({\rm GL}_{n-1}(L))\rightarrow\mathcal{R}(P_{n}(L))$
 $\Phi^{-}:\mathcal{R}(P_{n}(L))\rightarrow\mathcal{R}(P_{n-1}(L)),\Phi^{+}:\mathcal{R}(P_{n-1}(L))\rightarrow\mathcal{R}(P_{n}(L)),$
-given by $\Psi^{-}=r_{U_{n}(L),{\rm id}}$, $\Phi^{-}=r_{U_{n}(L),\theta_{n}}$, $\Psi^{+}=i_{U_{n}(L),{\rm id}}$, and $\Phi^{+}=i_{U_{n}(L),\theta_{n}}$.
+given by $\Psi^{-}=r_{U_{n}(L),{\rm id}}$, $\Phi^{-}=r_{U_{n}(L),\Theta_{L}}$, $\Psi^{+}=i_{U_{n}(L),{\rm id}}$, and $\Phi^{+}=i_{U_{n}(L),\Theta_{L}}$.
+6.3.1.
 Let $\tau$ be a smooth representation of $P_{n}(L)$.
@@
 $0\subseteq\tau_{n}\subseteq\tau_{n-1}\subseteq\dots\subseteq\tau_{2}\subseteq\tau_{1}=\tau,$
-where $\tau_{m}=(\Phi^{+})^{m-1}(\Phi^{-})^{m-1}(\tau)$ and $\tau_{m}/\tau_{m+1}=(\Phi^{+})^{m-1}(\Psi^{+})(\tau^{(m)})$. Using these, we now deduce:
+where $\tau_{m}=(\Phi^{+})^{m-1}(\Phi^{-})^{m-1}(\tau)$ and $\tau_{m}/\tau_{m+1}=(\Phi^{+})^{m-1}(\Psi^{+})(\tau^{(m)})$. We now deduce:
+Lemma 6.3.
+Let $\rho$ be a finite length representation of $G_{t}(L)$, where $1\leq t<n$. Then, the $P_{n}(L)$-representation $(\Phi^{+})^{n-t-1}(\Psi^{+})(\rho)$ is also of finite length.
+Proof.
+This is an immediate consequence of [Vig96, Chapter III, Subsection 1.5] and the exactness of the functor $(\Phi^{+})^{n-t-1}(\Psi^{+})$. ∎
+6.3.2.
+Let $F$ be a non-Archimedean local field with residue characteristic $p$.
+Let $E/F$ be a finite Galois extension of prime degree $l$ with $l\neq p$.
+Let $\sigma$ be a generator of ${\rm Gal}(E/F)$.
+Fix a non-trivial additive character $\psi_{F}:F\rightarrow k^{\times}$.
+Let $\psi_{E}$ be the character $\psi_{F}\circ{\rm Tr}_{E/F}$, where ${\rm Tr}_{E/F}$ is the trace function.
+We have the non-degenerate characters $\Theta_{F}$ and $\Theta_{E}$ of $N_{n}(F)$ and $N_{n}(E)$ respectively, defined by (6.1).
+We now prove the following finiteness result of Tate cohomology:
 Proposition 6.4.
@@
 Let $T_{\sigma}:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism with $T_{\sigma}^{l}={\rm id}$.
-Then the Tate cohomology $\widehat{H}^{i}(\Xi)$ with respect to the operator $T_{\sigma}$ is a finite length representation of ${\rm GL}_{n}(F)$.
+Then the Tate cohomology group $\widehat{H}^{i}(\Xi)$, with respect to the operator $T_{\sigma}$, has finite length as a representation of ${\rm GL}_{n}(F)$.
 Proof.
-We use induction on the integer $n$. Assume that the proposition is true for all positive integers less than $n$. Since $\psi_{E}(\sigma(x))=\psi_{E}(x)$, for $x\in E$, we get the isomorphism
+We prove the proposition using induction on $n$.
+The case $n=1$ is clear.
+So, we assume that the proposition is true for all irreducible $\sigma$-fixed representations of ${\rm GL}_{t}(E)$ and for all $t<n$.
+Now, we consider $\Xi$ as a representation of $P_{n}(E)$.
+Since $\psi_{E}(\sigma(x))=\psi_{E}(x)$, for $x\in E$, we get the isomorphism
 $\Phi^{-}(\Xi\circ\sigma)\simeq\Phi^{-}(\Xi)\circ\sigma,$ (6.2)
@@
 $\Phi^{+}(\tau\circ\sigma)\simeq\Phi^{+}(\tau)\circ\sigma.$ (6.3)
-The map $T_{\sigma}$ together with (6.1) and (6.2) induces an isomorphism between $\Xi^{(m)}$ and $(\Xi^{(m)})\circ\sigma$, and also between the representations $\Xi_{m}$ and $\Xi_{m}\circ\sigma$ for $m\geq 1$.
-Using induction hypothesis, we get that the representation $\widehat{H}^{i}(\Xi^{(m)})$ is a finite length representation of ${\rm GL}_{n-m}(F)$.
-Since $P_{n}(E)$ is stable under the action of $\sigma$, for any $n\geq 1$, we get that there is a bijection between the sets $X_{E}^{\sigma}$ and $X_{F}$.
-Then according to [TV16, Proposition 3.3.1], we have
-$\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})=(\Phi^{+})^{m-1}(\Psi^{+})\big{(}\widehat{H}^{i}(\Xi^{(m)})\big{)}.$
-Hence, for all $m\geq 1$, the representation $\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})$ is finite length representation of $P_{n}(F)$.
-Now, using spectral sequence induced by canonical filtration of ${\rm Res}_{P_{n}(E)}(\Xi)$, we get that $\widehat{H}^{i}(\Xi)$ is finite length representation of $P_{n}(F)$, and this implies that $\widehat{H}^{i}(\Xi)$ is finite length representation of ${\rm GL}_{n}(F)$.
+Using (6.2) and (6.3), and the isomorphism $T_{\sigma}$, we get an isomorphism between the representations $\Xi_{m}$ and $\Xi_{m}\circ\sigma$, and also between the representations $\Xi^{(m)}$ and $\Xi^{(m)}\circ\sigma$, for all $m\leq n$.
+Recall that $X_{E,m}$ denotes the coset space $P_{m}(E)/P_{m-1}(E)U_{m}(E)$.
+Since $P_{m-1}(E)U_{m}(E)$ is a $\sigma$-stable subgroup of $P_{m}(E)$, we have the following long exact sequence of non-abelian cohomology ([Ser, Appendix, Proposition 1]):
+$0\longrightarrow P_{m-1}(F)U_{m}(F)\longrightarrow P_{m}(F)\longrightarrow X_{E,m}^{\sigma}\longrightarrow H^{1}(\sigma,P_{m-1}(E)U_{m}(E))\longrightarrow H^{1}(\sigma,P_{m}(E))$ (6.4)
+Consider the short exact sequence of non-abelian ${\rm Gal}(E/F)$-modules
+$0\longrightarrow U_{m}(E)\longrightarrow P_{m-1}(E)U_{m}(E)\longrightarrow P_{m-1}(E)\longrightarrow 0$ (6.5)
+Using Hilbert’s theorem 90, we get that the pointed sets $H^{1}(\sigma,U_{m}(E))$ and $H^{1}(\sigma,P_{m-1}(E))$ are trivial.
+Then, it follows from the long exact sequence of non-abelian cohomology corresponding to (6.5) that $H^{1}(\sigma,P_{m-1}(E)U_{m}(E))=0$.
+Hence, the long exact sequence (6.4) gives the equality $X_{E,m}^{\sigma}=X_{F,m}$.
+Now, using Proposition 6.1 repeatedly $(m-1)$-times, we get the $P_{n}(F)$-equivariant isomorphism
+$\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})\simeq(\Phi^{+})^{m-1}(\Psi^{+})\big{(}\widehat{H}^{i}(\Xi^{(m)})\big{)}.$
+By induction hypothesis, the ${\rm GL}_{n-m}(F)$-representation $\widehat{H}^{i}(\Xi^{(m)})$ is of finite length, for all $m<n$.
+In view of Lemma 6.3, it then follows from the above isomorphism that the $P_{n}(F)$-representation $\widehat{H}^{i}(\Xi_{m}/\Xi_{m+1})$ is of finite length, for all $m<n$.
+Now, for each $m\in\{1,2,\dots,n-1\}$, consider the short exact sequence of $P_{n}(E)$-representations
+$0\longrightarrow\Xi_{m+1}\longrightarrow\Xi_{m}\longrightarrow\Xi_{m}/\Xi_{m+1}\longrightarrow 0$
+Since ${\rm Gal}(E/F)$ is cyclic, the corresponding long exact sequence of Tate cohomology gives the following diagram:
+${\widehat{H}^{0}(\Xi_{m+1})}$${\widehat{H}^{0}(\Xi_{m})}$${\widehat{H}^{1}(\Xi_{m}/\Xi_{m+1})}$${\widehat{H}^{0}(\Xi_{m}/\Xi_{m+1})}$${\widehat{H}^{1}(\Xi_{m})}$${\widehat{H}^{1}(\Xi_{m+1})}$
+We denote the above exact sequence by $S(m)$.
+Recall that $\Xi_{n}$ is the compactly induced representation ${\rm ind}_{N_{n}(E)}^{P_{n}(E)}(\Theta_{E})$, and using [TV16, Proposition 3.3.1], we get that $\widehat{H}^{0}(\Xi_{n})$ is isomorphic to the irreducible representation ${\rm ind}_{N_{n}(F)}^{P_{n}(F)}(\Theta_{F}^{l})$ and $\widehat{H}^{1}(\Xi_{n})=0$.
+In particular, the representation $\widehat{H}^{i}(\Xi_{n})$, for $i\in\{0,1\}$, is of finite length.
+Now, using the exact sequence $S(n-1)$ and the finiteness of $\widehat{H}^{i}(\Xi_{n-1}/\Xi_{n})$, we get that $\widehat{H}^{i}(\Xi_{n-1})$ is a finite length representation of $P_{n}(F)$.
+Again, using the finiteness of both the representations $\widehat{H}^{i}(\Xi_{n-1})$ and $\widehat{H}^{i}(\Xi_{n-2}/\Xi_{n-1})$, it follows from the exact sequence $S(n-2)$ that the representation $\widehat{H}^{i}(\Xi_{n-2})$ is of finite length.
+Thus, after a finite number of similar inductive steps, we finally get that $\widehat{H}^{i}(\Xi)$ is of finite length as a representation of $P_{n}(F)$, and hence of ${\rm GL}_{n}(F)$.
+This completes the proof.
 ∎
-It is reasonable to expect that Proposition 6.2 holds true for any general connected reductive algebraic group G over $F$.
-In the next section, we give a formulation of the compatibility of Brauer homomorphism with linkage, under the finite length assumption of the Tate cohomology $\widehat{H}^{i}(\Xi)$ for any irreducible $\sigma$-fixed representation $\Xi$ of $G=\textbf{G}(F)$.
-Then we show that this formulaton is compatible with Kazhdan isomorphism.
+It is reasonable to expect that Proposition 6.4 holds for arbitrary connected reductive algebraic groups G over $F$.
+Next, we give a formulation of the compatibility of Brauer homomorphism with linkage under the finite length assumption of the Tate cohomology $\widehat{H}^{i}(\Xi)$ for any irreducible $\sigma$-fixed representation $\Xi$ of $G=\textbf{G}(F)$.
+It enables us to study linkage from the module theoretic point of view via Brauer homomorphism.
+Then, using Theorem 5.1, we show that linkage is compatible with the Kazhdan isomorphism.
 6.4. Compatibility of Brauer homomorphism with linkage
 Let $\Xi$ be an irreducible $\sigma$-fixed representation of $G$, and let $T:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism such that $T^{l}={\rm id}$.
-For any $\sigma$-plain compact open subgroup $K$ of $G$, the inclusion map $\Xi^{K}\hookrightarrow\Xi$ is $\sigma$-invariant, and it induces a map $\widehat{H}^{i}(\Xi^{K})\rightarrow\widehat{H}^{i}(\Xi)$, taking values in the $\mathcal{H}(G^{\sigma},K^{\sigma})$-module $\widehat{H}^{i}(\Xi)^{K^{\sigma}}$.
-It is shown in [TV16, Section 6.2] that for all $h\in\mathcal{H}(G,K)^{\sigma}$, the Brauer homomorphism $\overline{\rm Br}$ induces the commutative diagram
+For any $\sigma$-plain compact open subgroup $K$ of $G$, the inclusion map $\Xi^{K}\hookrightarrow\Xi$ is $\sigma$-invariant and it induces a map $\widehat{H}^{i}(\Xi^{K})\rightarrow\widehat{H}^{i}(\Xi)$, taking values in the $\mathcal{H}(G^{\sigma},K^{\sigma})$-module $\widehat{H}^{i}(\Xi)^{K^{\sigma}}$.
+It is shown in [TV16, Section 6.2] that, for all $h\in\mathcal{H}(G,K)^{\sigma}$, the Brauer homomorphism $\overline{\rm Br}$ induces the following commutative diagram
 ${\widehat{H}^{i}(\Xi^{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{H}^{i}(h)}$${\widehat{H}^{i}(\Xi)^{K^{\sigma}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm\overline{Br}}(h)}$${\widehat{H}^{i}(\Xi^{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{i}(\Xi)^{K^{\sigma}}}$
-and this induces the compatibility of Brauer homomorphism with linkage in the sense that, for a representation $\rho$ of $G^{\sigma}$ to be linked with $\Xi$, it is equivalent to say that there exists a composition factor $N$ of the $\mathcal{H}(G,K)^{\sigma}$-module $\widehat{H}^{i}(\Xi^{K})$ such that
-$\mathcal{H}(G^{\sigma},K^{\sigma})\otimes_{\mathcal{H}(G,K)^{\sigma}}N\simeq(\rho^{(l)})^{K^{\sigma}}.$ (6.3)
+and it induces the compatibility of Brauer homomorphism with linkage in the sense that, for a representation $\rho$ of $G^{\sigma}$ to be linked with $\Xi$, it is equivalent to say that there exists a composition factor $\mathcal{N}$ of the $\mathcal{H}(G,K)^{\sigma}$-module $\widehat{H}^{i}(\Xi^{K})$ such that
+$\mathcal{H}(G^{\sigma},K^{\sigma})\otimes_{\mathcal{H}(G,K)^{\sigma}}\mathcal{N}\simeq(\rho^{(l)})^{K^{\sigma}}.$ (6.6)
 6.4.1. Compatibility of Kazhdan isomorphism with linkage
-We first recall the notations and set up under which we study this compatibility.
-Let $F$ and $F^{\prime}$ be the non-Archimedean $m$-close local fields with same residue characteristic $p$.
-Let $E$ be the finite Galois extension of $F$ of prime degree $l$ with $l\neq p$.
-By Lemma 4.2, there exists a finite Galois extension $E^{\prime}/F^{\prime}$ with same degree of extension $l$.
-As before, we fix the generators $\sigma$ and $\sigma^{\prime}$ of the Galois groups ${\rm Gal}(E/F)$ and ${\rm Gal}(E^{\prime}/F^{\prime})$ respectively.
+Let $F$ and $F^{\prime}$ be two non-Archimedean $m$-close local fields with the same residue characteristic $p$.
+Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$.
+By Lemma 4.6, we have a finite Galois extension $E^{\prime}/F^{\prime}$ of degree $l$ such that $E^{\prime}$ is $em$-close to $E$, where $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified.
+Fix two generators $\sigma$ and $\sigma^{\prime}$ of the groups ${\rm Gal}(E/F)$ and ${\rm Gal}(E^{\prime}/F^{\prime})$ respectively.
 Let G be a split connected reductive group defined over $\mathbb{Z}$.
-Let $G_{E}$ (resp.
-$G_{F}$) be the $E$ (resp.
-$F$)-points of G.
-Let $K_{E}$ and $K_{F}$ be the $m$-th congruence subgroups of $G_{E}$ and $G_{F}$ respectively (see Section 4).
-Then we have
-${\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}}={\rm\overline{Br}^{\prime}}\circ{\rm Kaz}_{em}^{E},$ (6.4)
-where $e\in\{1,l\}$. It is shown in [Gan15, Section 2.3] that there are bijections
-$\big{\{}\rho_{E}\in{\rm Irr}(G_{E}):\rho_{E}^{K_{E}}\neq 0\big{\}}\longrightarrow\big{\{}\rho_{E^{\prime}}\in{\rm Irr}(G_{E^{\prime}}):\rho_{E^{\prime}}^{K_{E^{\prime}}}\neq 0\big{\}}$ (6.5)
+Let $K_{E}$ (resp.
+$K_{F}$) be the congruence subgroup of $G_{E}$ (resp.
+$G_{F}$) of level $m$ (resp.
+$em$) (see Section 4 for the definition).
+Then, from Theorem 5.1, we get the following relation:
+${\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}}={\rm\overline{Br}^{\prime}}\circ{\rm Kaz}_{em}^{E}.$ (6.7)
+Moreover, we have the following bijections (see [Gan15, Section 2.3])
+$\big{\{}(\rho_{E},V_{E})\in{\rm Irr}(G_{E}):V_{E}^{K_{E}}\neq 0\big{\}}\longrightarrow\big{\{}(\rho_{E^{\prime}},V_{E^{\prime}})\in{\rm Irr}(G_{E^{\prime}}):V_{E^{\prime}}^{K_{E^{\prime}}}\neq 0\big{\}}$ (6.8)
 and
-$\big{\{}\rho_{F}\in{\rm Irr}(G_{F}):\rho_{F}^{K_{F}}\neq 0\big{\}}\longrightarrow\big{\{}\rho_{F^{\prime}}\in{\rm Irr}(G_{F^{\prime}}):\rho_{F^{\prime}}^{K_{F^{\prime}}}\neq 0\big{\}},$ (6.6)
-induced by the Kazhdan isomorphisms ${\rm Kaz}_{em}^{E}$ and ${\rm Kaz}_{m}^{F}$ respectively. With these notations and terminologies, we now prove
+$\big{\{}(\rho_{F},V_{F})\in{\rm Irr}(G_{F}):V_{F}^{K_{F}}\neq 0\big{\}}\longrightarrow\big{\{}(\rho_{F^{\prime}},V_{F^{\prime}})\in{\rm Irr}(G_{F^{\prime}}):V_{F^{\prime}}^{K_{F^{\prime}}}\neq 0\big{\}},$ (6.9)
+induced by the isomorphisms ${\rm Kaz}_{em}^{E}$ and ${\rm Kaz}_{m}^{F}$, respectively. With these notations, we now prove
 Theorem 6.5.
@@
 Let $(\rho_{E^{\prime}},V_{E^{\prime}})$ and $(\rho_{F^{\prime}},V_{F^{\prime}})$ be the corresponding objects under the bijections (6.8) and (6.9).
-Then $\rho_{F}$ is linked with $\rho_{E}$ if and only if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$.
+Then, the irreducible representation $\rho_{E^{\prime}}$ of $G_{E^{\prime}}$ is $\sigma^{\prime}$-fixed, and $\rho_{F}^{\prime}$ is linked with $\rho_{E}^{\prime}$ if and only if $\rho_{F}$ is linked with $\rho_{E}$.
 Proof.
-The proof is divided into two parts.
-In the first part, we show that the irreducible representation $\rho_{E^{\prime}}$, obtained from $\rho_{E}$ under the bijection (6.5), is $\sigma^{\prime}$-fixed.
-The second part shows that the linkage between $\rho_{F}$ and $\rho_{E}$ implies the linkage between $\rho_{F^{\prime}}$ and $\rho_{E^{\prime}}$, and vice versa.
-6.3.2.
-First recall that the space $V_{E}^{K_{E}}$ is an irreducible $\mathcal{H}(G_{E},K_{E})$-module, and it can be regarded as $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$-module via Kazhdan isomorphism.
-Consider the $\mathcal{H}(G_{E^{\prime}})$ (For definition, see [Vig96, Chapter 1])-module
-$N^{\prime}=\mathcal{H}(G_{E^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})}V_{E}^{K_{E}}.$
-Note that the action of $\langle\sigma^{\prime}\rangle$ on $\mathcal{H}(G_{E^{\prime}})$ induces an action of $\langle\sigma^{\prime}\rangle$ on $N^{\prime}$, given by
+We divide the proof into two parts.
+The first part proves that the irreducible representation $\rho_{E^{\prime}}$ is $\sigma^{\prime}$-fixed.
+In the second part, we show that the linkage between $\rho_{F}$ and $\rho_{E}$ implies the linkage between $\rho_{F^{\prime}}$ and $\rho_{E^{\prime}}$, and vice versa.
+6.4.2.
+Recall that the space of $K_{E}$-fixed vectors $V_{E}^{K_{E}}$ is a simple $\mathcal{H}(G_{E},K_{E})$-module, and it can be regarded as a $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$-module via the Kazhdan isomorphism ${\rm Kaz}_{em}^{E}$.
+Consider the $\mathcal{H}(G_{E^{\prime}})$-module
+$\mathcal{N}^{\prime}=\mathcal{H}(G_{E^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})}V_{E}^{K_{E}}.$
+Note that the action of $\langle\sigma^{\prime}\rangle$ on $\mathcal{H}(G_{E^{\prime}})$ induces an action of $\langle\sigma^{\prime}\rangle$ on $\mathcal{N}^{\prime}$, by setting
 $\sigma^{\prime}.(f^{\prime}\otimes v)=(\sigma^{\prime}.f^{\prime})\otimes v,$
-for $f^{\prime}\in\mathcal{H}(G_{E^{\prime}})$ and $v\in V_{E}^{k_{E}}$.
-Following the construction of [Vig96, Section 4, Chapter 1], the $\mathcal{H}(G_{E^{\prime}})$-module $N^{\prime}$ gives the irreducible representation $\rho_{E^{\prime}}$, defined as
+for all $f^{\prime}\in\mathcal{H}(G_{E^{\prime}})$ and $v\in V_{E}^{K_{E}}$.
+Following the construction of [Vig96, Chapter I, Subsection 4.4], the $\mathcal{H}(G_{E^{\prime}})$-module $\mathcal{N}^{\prime}$ gives the irreducible representation $(\rho_{E^{\prime}},V_{E^{\prime}})$, defined as
 $\rho_{E^{\prime}}(g^{\prime})(x^{\prime})=1_{g^{\prime}K_{E^{\prime}}}*x^{\prime},$
 for $g^{\prime}\in G_{E^{\prime}},x^{\prime}\in V_{E^{\prime}}$.
-Here, *** denotes the action of the Hecke algebra $\mathcal{H}(G_{E^{\prime}})$ on the space $V_{E^{\prime}}$ and $1_{g^{\prime}K_{E^{\prime}}}$ is the characteristic function of the left coset $g^{\prime}K_{E^{\prime}}$.
+Here * denotes the action of the Hecke algebra $\mathcal{H}(G_{E^{\prime}})$ on the space $V_{E^{\prime}}$ and $1_{g^{\prime}K_{E^{\prime}}}$ is the characteristic function of the left coset $g^{\prime}K_{E^{\prime}}$.
 Since $\rho_{E}$ is $\sigma$-fixed, there exists an isomorphism $T_{\sigma}:\rho_{E}\rightarrow\rho_{E}\circ\sigma$.
-Then the map $T_{\sigma^{\prime}}:N^{\prime}\rightarrow N^{\prime}$, defined by
+Then the map $T_{\sigma^{\prime}}:\mathcal{N}^{\prime}\rightarrow\mathcal{N}^{\prime}$, defined by
 $T_{\sigma^{\prime}}(f^{\prime}\otimes v):=(\sigma^{\prime}.f^{\prime})\otimes T_{\sigma}(v),$
-induces an isomorphism $\rho_{E^{\prime}}\xrightarrow{\sim}\rho_{E^{\prime}}\circ\sigma^{\prime}$, and this shows that the representation $\rho_{E^{\prime}}$ is $\sigma^{\prime}$-fixed.
-6.3.3.
-Now, suppose that the representation $\rho_{F}$ is linked with $\rho_{E}$.
-Then the $\mathcal{H}(G_{F},K_{F})$-module $(V_{F}^{(l)})^{K_{F}}$ is a composition factor of $\widehat{H}^{i}(V_{E})^{K_{F}}$, for some $i\in\{0,1\}$.
-If we regard these as $\mathcal{H}(G_{E},K_{E})^{\sigma}$-modules via the Brauer homomorphism $\overline{\rm Br}$, then the formulation (6.3) gives a composition factor $M$ of $\widehat{H}^{i}(V_{E}^{K_{E}})$ such that
-$\mathcal{H}(G_{F},K_{F})\otimes_{\mathcal{H}(G_{E},K_{E})^{\sigma}}M\simeq(V_{F}^{(l)})^{K_{F}}.$ (6.7)
+induces an isomorphism between $\rho_{E^{\prime}}$ and $\rho_{E^{\prime}}\circ\sigma^{\prime}$. Hence, $\rho_{E^{\prime}}$ is $\sigma^{\prime}$-fixed.
+6.4.3.
+Suppose the representation $\rho_{F}$ is linked with $\rho_{E}$.
+Then, the space of $K_{F}$-fixed vectors $(V_{F}^{(l)})^{K_{F}}$, considered as $\mathcal{H}(G_{F},K_{F})$-module, is a composition factor of $\widehat{H}^{i}(V_{E})^{K_{F}}$, for some $i$.
+If we regard these as $\mathcal{H}(G_{E},K_{E})^{\sigma}$-modules via the Brauer homomorphism $\overline{\rm Br}$, then the formulation (6.6) gives a composition factor $\mathcal{M}$ of $\widehat{H}^{i}(V_{E}^{K_{E}})$ such that
+$\mathcal{H}(G_{F},K_{F})\otimes_{\mathcal{H}(G_{E},K_{E})^{\sigma}}\mathcal{M}\simeq(V_{F}^{(l)})^{K_{F}}.$ (6.10)
 Note that the isomorphism $V_{E}^{K_{E}}\rightarrow V_{E^{\prime}}^{K_{E^{\prime}}}$ of $\mathcal{H}(G_{E},K_{E})$-modules induces a $\mathcal{H}(G_{E},K_{E})^{\sigma}$-module isomorphism $\Phi:\widehat{H}^{i}(V_{E}^{K_{E}})\rightarrow\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, which is also a $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}$-module isomorphism via the Kazhdan map ${\rm Kaz}^{E}_{em}$.
-Let $M^{\prime}$ be the image of $M$ under the map $\Phi$.
-Then $M^{\prime}$ is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, and the $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$-module
-$\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}M^{\prime}$
+Let $\mathcal{M}^{\prime}$ be the image of $\mathcal{M}$ under $\Phi$.
+Then $\mathcal{M}^{\prime}$ is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, and the $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$-module
+$\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}\mathcal{M}^{\prime}$
 is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}})^{K_{F^{\prime}}}$.
-We also have an isomorphism $\varphi:V_{F}^{K_{F}}\rightarrow V_{F^{\prime}}^{K_{F^{\prime}}}$, which is equivariant under the action of $\mathcal{H}(G_{F},K_{F})$ and hence $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$ via the Kazhdan isomorphism ${\rm Kaz}_{m}^{F}$.
+We also have an isomorphism $\varphi:V_{F}^{K_{F}}\rightarrow V_{F^{\prime}}^{K_{F^{\prime}}}$, which is equivariant under the action of $\mathcal{H}(G_{F},K_{F})$ and hence of $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$ via Kazhdan isomorphism ${\rm Kaz}_{m}^{F}$.
 Using the isomorphisms $\Phi$, $\varphi$ and the relation (6.7), it follows from (6.10) that
-$\mathcal{H}(G_{F,},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}M^{\prime}\simeq(V_{F^{\prime}}^{(l)})^{K_{F^{\prime}}}.$
-This shows that the representation $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$.
-A similar argument also establishes the other implication, namely that the representation $\rho_{F}$ is linked with $\rho_{E}$ if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$.
-This completes the proof.
+$\mathcal{H}(G_{F,},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}\mathcal{M}^{\prime}\simeq(V_{F^{\prime}}^{(l)})^{K_{F^{\prime}}}.$
+This implies that the representation $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$.
+A similar argument also shows that the representation $\rho_{F}$ is linked with $\rho_{E}$ if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$.
+Hence, the theorem.
 ∎
-Acknowledgements. The author is grateful to Santosh Nadimpalli for communicating Proposition 6.2, and for his helpful comments during the work.
+Acknowledgements.
+The author expresses his deep gratitude to the anonymous referee for careful reading of the manuscript and for various suggestions.
+The author is grateful to Santosh Nadimpalli for communicating Proposition 6.4, and for his helpful comments during the work.
+The author would also like to thank Radhika Ganapathy for her valuable comments and suggestions.